### v1

# DGC-GNN: Descriptor-free Geometric-Color Graph Neural Network for 2D-3D Matching

Abstract
Direct matching of 2D keypoints in an input image to a 3D point cloud of the scene without requiring visual descriptors has garnered increased interest due to its lower memory requirements, inherent privacy preservation, and reduced need for expensive 3D model maintenance compared to visual descriptor-based methods. However, existing algorithms often compromise on performance, resulting in a significant deterioration compared to their descriptor-based counterparts. In this paper, we introduce DGC-GNN, a novel algorithm that employs a global-to-local Graph Neural Network (GNN) that progressively exploits geometric and color cues to represent keypoints, thereby improving matching robustness. Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming the geometric embedding to guide the local point matching. We evaluate DGC-GNN on both indoor and outdoor datasets, demonstrating that it not only doubles the accuracy of the state-of-the-art descriptor-free algorithm but, also, substantially narrows the performance gap between descriptor-based and descriptor-free methods. The code and trained models will be made publicly available.
1 Introduction
Establishing 2D-3D matches plays a crucial role in various computer vision applications, including visual localization irschara2009structure ; sattler2012improving ; sattler2018benchmarking ; sarlin2019coarse ; wang2021continual ; sarlin2020superglue , 3D reconstruction snavely2006photo ; cui2015global ; schoenberger2016sfm ; lindenberger2021pixel , and Simultaneous Localization and Mapping (SLAM) eade2006scalable ; mur2015orb ; mur2017orb . Traditional methods for establishing point-to-point matches involve extracting keypoints and descriptors from a query image, then matching the 2D and 3D descriptors using exhaustive search. To circumvent the computationally expensive matching process, some approaches irschara2009structure ; sarlin2019coarse narrow the search space by employing image retrieval methods nister2006scalable ; Arandjelovic16 first to identify the most similar images in the database, and then perform descriptor-based image matching lowe2004distinctive ; detone2018superpoint ; dusmanu2019d2 ; sarlin2020superglue ; shi2022clustergnn between the query and retrieved images. The 2D-3D correspondences are subsequently established by connecting the 2D-2D image matches with the prebuilt 2D-3D correspondences in the database. Another approach sattler2011fast proposes an efficient vocabulary-based method to directly build 2D-to-3D matches by searching through all point descriptors. Sattler et al. sattler2012improving ; sattler2016efficient further explore the combination of both 2D-3D and 3D-2D search as an active correspondence search step for a faster and more efficient matching process.
While descriptor-based algorithms achieve state-of-the-art accuracy, they store and maintain high-dimensional visual descriptors for each point in potentially large 3D point clouds, significantly increasing both memory footprint and maintenance costs. The stored model often requires orders of magnitude more storage than the point cloud and images alone zhou2022geometry . These methods are susceptible to privacy attacks dosovitskiy2016inverting ; dosovitskiy2016generating ; chelani2021privacy and necessitate computationally expensive model maintenance and descriptor update procedures zhou2022geometry when incorporating new descriptors or points into the model. Several approaches have been proposed to address these limitations. Yang et al.Yang_2022_CVPR employ a learned point selection module to sample a subset of the point cloud for scene compression. Other methods li2020hierarchical ; brachmann2021dsacstar directly learn a function that maps 2D pixels to 3D coordinates without explicitly storing the 3D scene. Additionally, ng2022ninjadesc introduces an adversarial learning framework to develop content-concealing visual descriptors that prevent privacy leakage.
Recently, researchers campbell2020solving ; liu2020learning have begun exploring deep learning techniques for cross-domain direct 2D-3D matching and pose estimation without visual descriptors, showcasing the potential of descriptor-free matching through differentiable geometric optimization. The recently proposed GoMatch zhou2022geometry represents significant progress in descriptor-free 2D-3D matching, achieving reasonable matching performance on a variety of real-world datasets li2018megadepth ; shotton2013scene ; kendall2015posenet . GoMatch first identifies keypoints in the query image, which, along with the 3D points from the model, are converted to bearing vectors in the camera coordinate system. The algorithm employs the attention mechanism sarlin2020superglue ; vaswani2017attention to effectively establish reliable 2D-3D correspondences. While GoMatch attains reasonable accuracy, its performance still significantly lags behind its descriptor-based counterparts sarlin2020superglue ; sarlin2019coarse ; sattler2016efficient . Additionally, our experiments reveal that its reliance on geometric cues from the points and their local neighbors renders GoMatch incapable of distinguishing geometrically indistinct structures.
Figure 1: 2D-3D matching (shown by green lines) with the proposed DGC-GNN and GoMatch zhou2022geometry . In this example, DGC-GNN obtains 78 correct matches with 0.02 meters camera translation and ${}^{\circ}$ rotation errors, while GoMatch finds only 17 inliers with a pose error of 0.37 meters and ${}^{\circ}$.
These observations lead us to two critical questions: (1) Is geometry the only information we can utilize? (2) How can we leverage the geometric information derived from the points for matching? In practice, humans identify correspondences between objects by considering global structures and local geometric cues. For example, when matching an image to a point cloud as in Fig. 1, we first locate the building based on its unique structure and then identify the local structure of the roof for matching. Besides geometric cues, the visual context, such as the color information at each point, also provides constraints for 2D-3D matching. Importantly, this color information still preserves privacy, as the RGB data from sparse keypoints is insufficient to reconstruct the scene.
Building upon these observations and the groundwork set by GoMatch, we propose a novel graph-based pipeline, named DGC-GNN, which leverages geometric and color cues in a global-to-local manner for descriptor-free 2D-3D matching. Our pipeline encodes both position and RGB information for each point and extracts a global distance-angular embedding to guide local point matching. Taking inspiration from shi2022clustergnn , we employ a cluster-based transformer to constrain information flow within local clusters. We observe, from real-world datasets, that the proposed DGC-GNN leads to substantial improvements in the number of correct matches and the accuracy of pose estimation. Notably, it doubles the accuracy of GoMatch, thereby reducing the gap between descriptor-based and descriptor-free methods. In summary, our paper makes the following contributions:
  • We introduce a descriptor-free global-to-local GNN for direct 2D-3D point matching. The network leverages multiple cues and incorporates a progressive clustering module to represent the keypoints. This pipeline enhances the robustness and accuracy of 2D-3D matching while requiring low memory, being privacy-preserving and free from 3D model maintenance.
  • We demonstrate that color information is crucial for 2D-3D matching. By incorporating RGB encoding into our network, we observe significant improvements in performance.
  • Extensive experiments on multiple real-world datasets show that the proposed network outperforms previous methods by a large margin on both matching and visual localization tasks.
2 Descriptor-Free 2D-3D Matching
2.1 Problem Formulation and Notation
Considering keypoints $\mathbf{P}=\{\mathbf{p}_{n}\in\mathbb{R}^{2}\;|\;n=1,...,N\}$ from a query image $I$ and a database 3D point cloud $\mathbf{Q}=\{\mathbf{q}_{m}\in\mathbb{R}^{3}\;|\;m=1,...,M\}$, where, optionally, each 3D point is associated with a visual descriptor $\mathbf{d}\in\mathbb{R}^{D}$. The task of 2D-3D point matching is to find a set $\mathcal{M}_{\mathbf{p,q}}$ of corresponding keypoints such that
$\mathcal{M}_{\mathbf{p,q}}=\{(n,m)\;|\;||\pi(\mathbf{q}_{m},\mathbf{R,t,K})-\mathbf{p}_{n}||_{2}\leq\epsilon\},$ (1)
where $\pi(\cdot)$ is a mapping that projects a 3D point $\mathbf{q}_{m}$ from world coordinates to the image plane, represented by a camera rotation $\mathbf{R}\in\mathbb{R}^{3\times 3}$, translation $\mathbf{t}\in\mathbb{R}^{3}$, and intrinsic parameter matrix $\mathbf{K}\in\mathbb{R}^{3\times 3}$. Parameter $\epsilon\in\mathbb{R}$ is the threshold specified in pixels. Additionally, we denote the color of point $\mathbf{p}_{n}$ as $\mathbf{c}_{n}=[r,g,b]^{\text{T}}\in[0,1]^{3}$.
Bearing Vector. Similar to zhou2022geometry , we adopt bearing vectors as keypoint representation for both the 2D and 3D points to alleviate their cross-domain nature and represent them in the same space. The bearing vector is the direction from the camera center to a 3D point in the camera coordinate system. Given an image, a 2D pixel $\mathbf{p}_{n}$ is uplifted to bearing vector as $[\mathbf{b}_{\mathbf{p},n},1]^{\text{T}}=\mathbf{K^{-1}}[\mathbf{p}_{n},1]^{\text{T}},\mathbf{b}_{\mathbf{p},,}\in\mathbb{R}^{2}$, where $\mathbf{K}$ is the intrinsic camera matrix. Given a 3D point $\mathbf{q}_{m}$, the corresponding bearing vector is $[\mathbf{b}_{\mathbf{p},m},1]^{\text{T}}=\frac{\mathbf{Rq}_{m}+\mathbf{t}}{[\mathbf{Rq}_{m}+\mathbf{t}]_{z}}$, where $\mathbf{R}$ is the camera rotation and $\mathbf{t}$ is its translation in the world coordinate system and subscript $z$ denotes the third component of the 3D vector.
2.2 Network Architecture
Figure 2: Pipeline overview. For keypoints from the 2D image and 3D points from the point cloud, the proposed DGC-GNN (1) considers the bearing vectors and the color at each bearing vector as input. (2) It extracts the point-wise position and color features with two separate encoders and mixes the features as $\mathbf{f_{p}}$ and $\mathbf{f_{p}}$. (3) The bearing vectors are clustered into $K$ groups, and geometric graphs are built upon the clusters to extract the global-level geometric embeddings $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$. (4) We then concatenate $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ with $\mathbf{f_{p}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$ with $\mathbf{f_{p}}$, and build a local graph at each point as self-attention. A cluster-based attention module is adopted to enhance the local features by forcing the message passing only with the most related features. A differentiable layer matches and optimizes the improved features to obtain score matrix $\mathcal{S}$. Finally, an outlier filtering network is applied to prune the matches with low confidence, leading to the final 2D-3D correspondences $\mathcal{M}_{final}$.
The proposed DGC-GNN applies a hierarchical mechanism to leverage both color and geometric cues in a global-to-local fashion. The overall pipeline is illustrated in  Fig.2. We initially employ two local feature extractors to encode RGB and position information for each point simultaneously (Sec. 2.2.1). Additionally, we cluster the points based on their distances and generate global graphs to obtain the global-level geometric embeddings (Sec. 2.2.2). Next, we concatenate the local point features with their corresponding global features and input them into the cluster-based local matching module to identify the initial matches (Sec. 2.2.3). Finally, we incorporate a classification network to filter out matches with low confidence to refine the initial matches (Sec.2.3).
2.2.1 Local Feature Extraction
To extract points-wise features from both the 2D keypoint set $\mathbf{P}$ and the 3D point cloud $\mathbf{Q}$, we consider the inputs as bearing vectors equipped with color information: $\mathcal{P}=\{\mathbf{b_{p},c_{p}}\}$ and $\mathcal{Q}=\{\mathbf{b_{q},c_{q}}\}$. Two ResNet-style point encoders he2016deep ; campbell2020solving , denoted as $\mathcal{F}_{b}$ and $\mathcal{F}_{c}$, are applied to extract position and color embeddings separately. We then obtain the local point features, $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$, as follows:
$\mathbf{f_{p}}=\mathcal{F}_{b}(\mathbf{b_{p}})+\mathcal{F}_{c}(\mathbf{c_{p}}),\ \ \mathbf{f_{q}}=\mathcal{F}_{b}(\mathbf{b_{q}})+\mathcal{F}_{c}(\mathbf{c_{q}}).$ (2)
The resulting point-wise features $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$ are vectors with dimensions $\mathbb{R}^{N\times d}$ and $\mathbb{R}^{M\times d}$ respectively, where $N$ and $M$ represent the number of keypoints in $\mathbf{P}$ and $\mathbf{Q}$, and $d$ denotes the dimensionality of the encoded features, e.g., $d=128$ for SIFT features lowe2004distinctive .
2.2.2 Global Geometric Guidance
Global context guidance has demonstrated its effectiveness in various computer vision tasks sun2021loftr ; li2020hierarchical ; yu2021cofinet ; qin2022geometric . Global context helps to differentiate local descriptors from similar structures or patches, thereby reducing ambiguity. However, most existing methods sun2021loftr ; qin2022geometric consider the outputs from different encoding layers as global and local features. This approach is not suitable for our scenario, as our input is sparse points. Downsampling the sparse point cloud results in losing distinctive geometric structures. Hence, we adopt cluster-based geometric encoding to extract global embeddings. As shown in Fig. 3.(a) and (c), the input bearing vectors, both in the image and in the point cloud, are first clustered into $K$ groups. The groups represent distinct clusters, each associated with a cluster center as the global position, denoted by $\mathbf{\hat{b}}_{\mathbf{p},k}\in\mathbb{R}^{2},k=1,...K$. The corresponding global embedding is obtained by computing the average of the point embeddings within the cluster, thus $\mathbf{\hat{f}}_{\mathbf{p},k}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{f}_{\mathbf{p},k},\ \mathbf{\hat{f}}_{\mathbf{p},k}\in\mathbb{R}^{d}$. The same is conducted on the 3D points to obtain $\mathbf{\hat{b}}_{\mathbf{q}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}$.
Figure 3: Cluster-based geometric encoding. (a) The clusters obtained from bearing vectors $\mathcal{Q}$ of the 3D point cloud are visualized by color. The local graph is created from the neighboring cluster centers. Black 3D points are filtered out from matching. (b) Angular embedding from the global graph to obtain rotation-invariant geometric cues. (c) The clusters obtained from 2D keypoints’ bearing vectors $\mathcal{P}$. Similarly, as in 3D, the local graph is created from the neighboring cluster centers.
Global Geometric Graph. To aggregate and extract the geometric relations among the clusters, we propose a novel graph neural network that encodes both distance and angular cues; the basic GNN structure is built upon huang2021predator ; zhou2022geometry . In the following, we describe the graph building for the 2D global points set $\mathcal{\hat{P}}=\{\mathbf{\hat{b}_{p}},\mathbf{\hat{f}_{p}}\}$ and the same goes for $\mathcal{\hat{Q}}=\{\mathbf{\hat{b}_{q}},\mathbf{\hat{f}_{q}}\}$. Each cluster center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ is connected to its $k$-NN neighbours ($k\leq K$) in the coordinate space, and $\xi_{\mathbf{p},(x,y)}$ is the edge between center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$. We update the feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ using the following equation:
${}^{(t+1)}\mathbf{\hat{f}}_{\mathbf{p},x}=\max_{\xi_{\mathbf{p},(x,y)}}\mathcal{H}_{g}(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}-^{(t)}\mathbf{\hat{f}}_{\mathbf{p},y})),$ (3)
where the $\oplus$ denotes concatenation and $\mathcal{H}_{g}(*)$ is the linear projection with instance normalization ulyanov2016instance and a LeakyReLU function maas2013rectifier . The global feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ is updated twice, and calculated as
$\mathbf{\hat{f}}_{\mathbf{p},x}^{g}=\mathcal{H}_{g}(^{(0)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(1)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(2)}\mathbf{\hat{f}}_{\mathbf{p},x}).$ (4)
Besides the distance embedding, inspired by qin2022geometric , we also adopt the angular embedding to obtain rotation-invariant geometric cues for the global point representation. To do so, we define the embedding on cluster triplets as shown in Fig. 3.(b). Given bearing vector $\mathbf{\hat{b}}_{\mathbf{p},z}$ and two of its neighbors $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$, the angular embedding of $\mathbf{\hat{b}}_{\mathbf{p},z}$ is defined as follows:
$\mathbf{A}_{x,y}^{z}=\operatorname{sine}(\angle(\mathbf{\hat{b}}_{\mathbf{p},z}-\mathbf{\hat{b}}_{\mathbf{p},x},\ \ \ \mathbf{\hat{b}}_{\mathbf{p},y}-\mathbf{\hat{b}}_{\mathbf{p},x})/\sigma_{a}),$ (5)
where $\operatorname{sine}(\cdot)$ is a sinusoidal function and $\sigma_{a}$ is a controller constant. We update the global geometric embedding $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ as an angular-aware attention mechanism as follows:
$\mathbf{\hat{f}}_{\mathbf{p}}^{gg}=\operatorname{norm}(\mathbf{\hat{f}}_{\mathbf{p}}^{g}+\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}}));\ \ \ \ \mathbf{\hat{f}}_{\mathbf{p}}^{gg}\in\mathbb{R}^{K\times d},$ (6)
where
$\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}})=(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{V}}).\frac{(\mathbf{A}_{\mathbf{p}}\mathbf{W^{A}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})^{\text{T}}+(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{K}})^{\text{T}}}{\sqrt{dim}}.$ (7)
$\mathbf{W^{A}},\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{d\times d}$ are the projection matrices of each item. The local point representation with global geometric embedding is as $\mathcal{\widetilde{P}}=\{\mathbf{b}_{\mathbf{p}},\mathbf{\widetilde{f}}_{\mathbf{p}}\}$ and $\mathbf{\widetilde{f}}_{\mathbf{p}}=\mathbf{f}_{\mathbf{p}}\oplus\mathbf{\hat{f}}_{\mathbf{p}}^{gg},\mathbf{\widetilde{f}}_{\mathbf{p}}\in\mathbb{R}^{N\times 2d}$. The operation for the bearing vectors obtained from point cloud $\mathcal{\hat{Q}}$ is the same and we get $\mathcal{\widetilde{Q}}=\{\mathbf{b}_{\mathbf{q}},\mathbf{\widetilde{f}}_{\mathbf{q}}\}$.
2.2.3 Cluster-based Local Matching
After extracting the global geometric embedding, we implement a cluster-based matching module to obtain the initial intra-domain 2D-3D matches. This cluster-based GNN  shi2022clustergnn has been shown to be more computationally efficient than its complete-graph counterpart sarlin2020superglue . The network considers the local point features from both $\mathcal{\widetilde{P}}$ and $\mathcal{\widetilde{Q}}$ a complete set, then classifies the local feature with strong correlations into the same group and restricts the message passing within each group. In addition to its low computational complexity, we found that the cluster GNN can effectively utilize our global-to-local geometric cues, as the clustering operation inherits the property of global graph clustering and forces it to distinguish ambiguous local features even with similar global embedding.
GoMatch-style Initialization. Our local graph initialization follows the pipeline in GoMatch zhou2022geometry and conducts it as geometric self-attention. For each local point $\mathbf{b}_{\mathbf{p},n}$, we construct a local graph according to its $k^{\prime}$ nearest neighbours in the Euclidean space and update the associated feature $\mathbf{\widetilde{f}}_{\mathbf{p},n}\in\mathbb{R}^{2d}$ by Eq. 4. Note that we ignore the angular embedding at this stage due to the unaffordable memory requirements with space complexity $\mathcal{O}(Nk^{\prime 2})$, where $N$ is the number of local points. We then adopt linear attention katharopoulos2020transformers ; sun2021loftr as a cross-attention mechanism, which allows each keypoint in one modality to interact with all keypoints from another modality. This not only facilitates inter-modality in the feature matching but also reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
Cluster-based Attention. After the graph initialization, the features $\mathbf{\widetilde{f}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}}_{\mathbf{q}}$ coming from the image and the point cloud respectively, are concatenated and processed in a two-level hierarchical clustering attention module. The hierarchical structure is effective in reducing erroneous grouping. At the first level, we directly cluster the feature vectors into $I$ coarse groups. In the second level, each coarse group is divided into several small groups. The local point information exchange is conducted at each level and only within each group to obtain more representative features. After the sparse clustering, each feature vector is transformed back to its original position and then split again into $\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q}}$ to obtain the keypoints both in the 2D and 3D spaces.
[…]
2.3 Outlier Rejection
After obtaining the initial matches, an outlier pruning approach is adopted to remove the incorrect ones. We apply a classification network, as applied in yi2018learning ; zhou2022geometry , whose input is the concatenated 2D and 3D keypoint features $\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}=\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p},\widetilde{n}}\oplus\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q},\widetilde{m}}$ and outputs the matching confidence of each matched pair. The final predicted matches are obtained as follows:
$\mathcal{M}_{final}=\{(\widetilde{n}^{\prime},\widetilde{m}^{\prime})\ |\ \forall\ \operatorname{classifier}(\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}\;|\;(\widetilde{n},\widetilde{m})\in\mathcal{M}_{init})\geq\theta\},$ (8)
where $\theta$ is the matching confidence threshold.
2.4 Training Loss
We use the same training loss as GoMatch. The loss function $\mathcal{L}$ consists of two terms, the matching loss $\mathcal{L}_{ot}$ and the classification loss $\mathcal{L}_{or}$. The ground truth match set $\mathcal{M}_{gt}$ is estimated by reprojecting the 3D points to the 2D image plane and calculating the pixel distance. We also include point sets $\mathcal{I}$ and $\mathcal{J}$ for the unmatched points in $\mathcal{P}$ and $\mathcal{Q}$, respectively. The matching loss $\mathcal{L}_{ot}$ minimizes the negative log-likelihood of the matching score $\mathcal{\bar{S}}$.
$\mathcal{L}_{ot}=-\frac{1}{|\mathcal{M}_{gt}|+|\mathcal{I}|+|\mathcal{J}|}(\sum\limits_{(n,m)\in\mathcal{M}_{gt}}\log\mathcal{\bar{S}}_{n,m}+\sum\limits_{n\in\mathcal{I}}\log\mathcal{\bar{S}}_{i,m+1}+\sum\limits_{m\in\mathcal{J}}\log\mathcal{\bar{S}}_{N+1,j}).$ (9)
The classification loss is defined as
$\mathcal{L}_{or}=-\frac{1}{|\mathcal{M}_{init}|}\sum_{i=1}^{|\mathcal{M}_{init}|}w_{i}(y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})),$ (10)
where $w_{i}$ is the balance weight of positive and negative samples, $y_{i}$ is the gt matching label for the $i$-th correspondences, $p_{i}$ is the predicted probability of a true match for the $i$-th correspondences. The total loss is the sum of the two terms, thus, $\mathcal{L}=\mathcal{L}_{ot}+\mathcal{L}_{or}$.
3 Experiments
In this section, we first describe the experimental setup, employed datasets, and evaluation protocols. Next, we provide detailed comparisons with the baselines on both matching and visual localization tasks. Finally, we conduct ablation studies on each designed component.
Training. We train the indoor model of DGC-GNN on the ScanNet dai2017scannet dataset and the outdoor model on the MegaDepth li2018megadepth dataset. We extract up to 1024 keypoints for each training image by the SIFT detector lowe2004distinctive . Similarly as in GoMatch, we first select a subset of the point cloud by applying image retrieval approaches Arandjelovic16 ; torii201524 to obtain potential images observing the same part of the scene as the input one. During the training, we randomly sample the retrieval pairs with a visual overlap of more than 35% on MegaDepth and 65% on ScanNet to ensure enough matches on each pair. For the global geometric embedding, we cluster the 2D/3D bearing vectors into $K=10$ groups, and each cluster center point is connected to its $k=4$ nearest neighbors to build the global graph. For the local point graph, we connect each point with its 10 nearest neighbors and the cluster-based attentions are performed twice to force the intra-cluster information exchange.
The model is optimized using the Adam with a fixed learning rate of 1e-3. We train DGC-GNN with one 32GB Telsa V100 GPU. The convergence of the model typically requires 25 epochs with around 20 hours for indoor scenes and 35 epochs (approx. 35 hours) for outdoor scenes.
Datasets. We use ScanNet and MegaDepth for the model training and 2D-3D matching task evaluation. As a downstream application, we perform visual localization on the 7Scenes shotton2013scene and Camdridge-Landmarks kendall2015posenet datasets. MegaDepth is a popular outdoor dataset with 196 scenes captured around the world. The sparse 3D reconstructions are provided by the COLMAP schoenberger2016sfm structure-from-motion software. Following zhou2022geometry , we train our outdoor model on 99 scenes and evaluate it with 53 scenes. ScanNet is a large-scale RGB-D indoor dataset that consists of 1613 scans with over 2.5 million images. We randomly selected 105 scenes for the training and 30 scenes for the evaluation. Cambridge-Landmarks is a middle-scale outdoor dataset consisting of 6 individual scenes. The ground truth camera poses are provided by a structure-from-motion algorithm. We follow kendall2015posenet ; zhou2022geometry to evaluate our method on four of the scenes. 7scenes is a small indoor dataset with RGB-D images and camera poses provided by the depth SLAM system. We evaluate on the standard test sequences.
Evaluation Protocol. For matching on ScanNet and MegaDepth, we follow zhou2022geometry and report the AUC score calculated from the reprojection errors. To calculate the reprojection errors for the 2D-3D matches in $\mathcal{M}_{final}$, we project the 3D points to the image plane using the ground truth and estimated camera poses. Then we calculate the $L_{2}$ distance between the ground truth and estimated reprojected 2D points. We use multiple thresholds, 1, 5, and 10 pixels, to evaluate the algorithms. The camera translation and rotation error quantiles at 25%, 50%, and 75% are also reported. Besides, we also evaluate the matching quality by calculating the matching precision P, which is the ratio of inlier matches after PnP-RANSAC to the number of final matches $\mathcal{M}_{final}$. For visual localization tasks, we report the median translation (in meters) and rotation (in degrees) camera pose errors.
Methods ScanNet dai2017scannet MegaDepth li2018megadepth
Reproj. AUC (%) Rotation (${}^{\circ}$) Translation (m) P (%) ($\uparrow$) Reproj. AUC (%) Rotation (${}^{\circ}$) Translation P (%) ($\uparrow$)
@1 / 5 / 10px ($\uparrow$) Quantile @25 / 50 / 75% ()($\downarrow$)( ↓ ) @1 / 5 / 10px ($\uparrow$) Quantile @25 / 50 / 75% ()($\downarrow$)( ↓ )
k=1 Oracle 29.13 / 39.83 / 41.34 110.10 / 110.19 / 110.40 0.01 / 0.01 / 0.03 - 34.59 / 85.02 / 92.02 10.04 / 10.06 / 10.12 0.00 / 0.01 / 0.01 -
BPnPNet campbell2020solving 10.00 / 10.00 / 10.02 199.17 / 128.90 / 154.68 4.35 / 6.82 / 9.86 13.60 10.22 / 10.63 / 10.89 16.13 / 32.01 / 61.58 1.67 / 3.17 / 5.44 12.95
GoMatch zhou2022geometry 11.18 / 11.23 / 18.01 112.69 / 112.78 / 136.50 0.19 / 0.91 / 2.63 13.18 15.67 / 22.43 / 28.01 10.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73 14.94
DGC-GNN 12.73 / 21.88 / 32.23 110.94 / 113.17 / 120.14 0.06 / 0.23 / 1.40 14.86 10.20 / 37.64 / 44.04 10.15 / 11.53 / 27.93 0.01 / 0.15 / 3.00 19.00
k=10 BPnPNet campbell2020solving 10.00 / 10.00 / 10.03 104.68 / 135.94 / 160.54 4.67 / 7.30 / 10.92 10.84 10.36 / 10.72 / 10.97 16.63 / 34.69 / 67.77 1.64 / 3.30 / 5.97 10.74
GoMatch zhou2022geometry 10.91 / 18.98 / 31.12 111.18 / 114.94 / 128.97 10.08 / 0.35 / 2.08 14.25 18.90 / 35.67 / 44.99 10.18 / 11.29 / 16.65 0.02 / 0.12 / 1.92 18.76
DGC-GNN 11.76 / 31.74 / 48.11 110.67 / 111.49 / 117.62 10.04 / 0.11 / 0.53 16.42 15.30 / 51.70 / 60.01 10.07 / 10.26 / 15.41 0.01 / 0.02 / 0.57 13.36
Table 1: 2D-3D matching. We report the reprojection error AUC scores thresholded at 1, 5, and 10 pixels; rotation (degrees) and translation error (meters) quantiles at 25, 50, and 75%; and matching precision (percentages). On MegaDepth li2018megadepth , the translation scale is unknown, so the unit of the errors comes from the reference reconstruction and is not in meters. Parameter $k$ is the number of images retrieved from the database to narrow down the search space. The best results are bold. DGC-GNN nearly doubles the AUC scores of GoMatch and reduces the pose errors to their ${\approx}33\%$.
3.1 2D-3D Matching
We compare with the two descriptor-free matchers GoMatch zhou2022geometry and BPnPnet campbell2020solving . During inference, we use the 3D points from the top-$k$ retrieved database images to match against the keypoints from query images. Following zhou2022geometry , we report the upper bound of the AUC score at 1, 5, and 10 pixels by the ground truth matches. We refer to these values as Oracle. We use the official code with the default setting to generate the evaluation dataset on MegaDepth li2018megadepth and rerun the GoMatch and BPnPNet with the released models on the dataset. Note that we also tested GoMatch after retraining it on MegaDepth and achieved similar results as with the released model.
Matching Results. The results with $k=1$ and $k=10$ are presented in Table 1. $k$ is the number of retrieved image pairs that are used for evaluation. The proposed method outperforms GoMatch and BPnPNet by a significant margin on both scenes. Specifically, DGC-GNN achieves 10.2 / 37.64 / 44.04% reprojection AUC compared to GoMatch with 5.67 / 22.43 / 28.01% on MegaDepth with $k=1$. DGC-GNN halves the rotation and translation errors of GoMatch on all thresholds and it obtains better matching quality. Notably, the performance of DGC-GNN with $k=1$ surpasses that of GoMatch with $k=10$, indicating the effectiveness of our method even with a single view.
Figure 4: Outlier Sensitivity. The AUC scores of BPnPNet campbell2020solving , GoMatch zhou2022geometry and the proposed DGC-GNN thresholded at 1, 5, and 10 pixels are plotted as a function of the outlier ratio. Oracle represents the AUC score upper bound using ground truth 2D-3D matches.
Sensitivity to Outliers. To evaluate the sensitivity of our method to keypoint outliers, we follow the procedure in GoMatch zhou2022geometry . The outliers are controlled by the outlier ratio, ranging from 0 to 1, calculated as the number of unmatched keypoints divided by the maximum of the numbers of 2D and 3D points. In case the outlier ratio is 00, all of the input 2D and 3D points are selected from the ground truth matches, and no outliers are included in the matching process. When it is $1$, we directly use the keypoints from the 2D query and 3D points from top-$k$ retrieved images as input to find the matches without any filtering or outlier removal. The results are shown in Fig. 4. Even in the presence of outliers, DGC-GNN outperforms other methods by a large margin. This indicates that our method is more robust to the presence of outliers and can handle challenging matching scenarios more effectively than the state-of-the-art.
Methods G. Emb. C. Att. Color Ang. Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
Sec. 2.2.2 Sec. 2.2.3 Sec. 2.2.1 Sec. 2.2.2 @1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
GoMatch zhou2022geometry 18.90 / 35.67 / 44.99 0.18 / 1.29 / 16.65 0.02 / 0.12 / 1.92
Variants 10.86 / 41.18 / 50.51 0.13 / 0.76 / 13.47 0.01 / 0.07 / 1.62
11.64 / 44.46 / 53.99 0.11 / 0.55 / 19.49 0.01 / 0.05 / 1.05
14.19 / 48.34 / 56.54 0.08 / 0.34 / 19.23 0.01 / 0.03 / 1.03
DGC-GNN 15.30 / 51.70 / 60.01 0.07 / 0.26 / 15.41 0.01 / 0.02 / 0.57
Table 2: Ablation Study. AUC scores thresholded at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, 75% with the proposed components added one by one to the GoMatch pipeline.
Ablation Study. We investigate the effectiveness of different components of DGC-GNN on the 2D-3D matching quality. The impact of different components of DGC-GNN is studied on MegaDepth dataset li2018megadepth with $k=10$. The results are reported in Table 2. We provide other results in the supplementary material. We conduct the ablations by gradually adding the components: global geometric embedding (G. Emb), cluster attention (C. Att.), Color, and Angular embedding (Ang.) to the original GoMatch pipeline. Incorporating color information into the matching process significantly impacts the performance, resulting in improvements of 2.55 / 3.88 / 2.55% (AUC@1 / 5 / 10px). This demonstrates the importance of considering color cues for accurate and robust matching. The global-to-local geometric embedding (G. Emb.) and Angular relation embedding (Ang.) substantially improve performance by 1.90 / 5.51 / 5.52% and 1.11 / 3.36 / 3.47%, respectively. It highlights the effectiveness of our idea incorporating global geometric context and local geometric details. The cluster attention mechanism also plays a vital role, improving performance by 0.78 / 3.28 / 3.48%. The best results are obtained when all components are added to the pipeline.
Methods No Desc. Maint. Privacy Cambridge-Landmarks kendall2015posenet (cm, ${}^{\circ}$) MB used 7Scenes shotton2013scene (cm, ${}^{\circ}$) MB used
King’s Hospital Shop St. Mary’s Chess Fire Heads Office Pumpkin Kitchen Stairs
E2E MS-Trans. shavit2021learning 83 / 1.47 181 / 2.39 86 / 3.07 162 / 3.99 1171 11 / 4.66 24 / 9.60 14 / 12.19 17 / 5.66 18 / 4.44 17 / 5.94 26 / 8.45 11171
DSAC* brachmann2021dsacstar 15 / 0.30 121 / 0.40 15 / 0.30 113 / 0.40 1112 12 / 1.10 12 / 1.24 11 / 1.82 13 / 1.15 14 / 1.34 14 / 1.68 13 / 1.16 11196
HSCNet li2020hierarchical 18 / 0.30 119 / 0.30 16 / 0.30 119 / 0.30 1592 12 / 0.70 12 / 0.90 11 / 0.90 13 / 0.80 14 / 1.00 14 / 1.20 13 / 0.80 11036
DB HybridSC camposeco2019hybrid 81 / 0.59 175 / 1.01 19 / 0.54 150 / 0.49 1113 - - - - - - - -
AS sattler2016efficient 13 / 0.22 120 / 0.36 14 / 0.21 118 / 0.25 1813 13 / 0.87 12 / 1.01 11 / 0.82 14 / 1.15 17 / 1.69 15 / 1.72 4 / 1.01 -
SP detone2018superpoint +SG sarlin2020superglue 12 / 0.20 115 / 0.30 14 / 0.20 117 / 0.21 3215 12 / 0.85 12 / 0.94 11 / 0.75 13 / 0.92 15 / 1.30 14 / 1.40 15 / 1.47 22977
DF GoMatch zhou2022geometry 25 / 0.64 283 / 8.14 48 / 4.77 335 / 9.94 1148 14 / 1.65 13 / 3.86 19 / 5.17 11 / 2.48 16 / 3.32 13 / 2.84 89 / 21.12 11302
DGC-GNN 18 / 0.47 75 / 2.83 15 / 1.57 106 / 4.03 1169 13 / 1.43 15 / 1.77 14 / 2.95 16 / 1.61 18 / 1.93 18 / 2.09 71 / 19.5 11355
Table 3: Visual localization. We report the median pose errors (cm, ${}^{\circ}$) and storage requirements (MB) on the scenes of the 7Scenes shotton2013scene and Cambridge-Landmarks kendall2015posenet datasets. Three groups of methods are shown: end-to-end (E2E), descriptor-based (DB), and descriptor-free (DF). We do not show BPnPNet as it fails on most scenes. The best results are shown in bold in each group.
3.2 Visual Localization
The visual localization task is the estimation of the 6 degrees-of-freedom camera pose of an input query image w.r.t a known map of the scene. One of the most prominent ways of approaching this problem is via establishing 2D-3D correspondences and running robust pose estimation. Following zhou2022geometry , we ran the proposed DGC-GNN to obtain matches. For each query image, we match its keypoints against the 3D points from the top-10 retrieved views to build the 2D-3D correspondences. The camera pose is then estimated by PnP-RANSAC fischler1981random ; 1217599 . We consider two standard datasets 7scenes shotton2013scene and Cambridge landmarks kendall2015posenet to evaluate the proposed method. For 7scenes, we extract the keypoints with the SIFT detectors, and the top 10 pairs are retrieved using DenseVLAD torii201524 . For Cambridge landmarks, the keypoints are extracted by SuperPoint detone2018superpoint to ensure consistency with the SuperPoint-based structure-from-motion model. The top 10 pairs are provided by NetVLAD Arandjelovic16 .
Results. In Table 3, we present the 3D model maintenance costs, privacy, storage requirements, and camera pose median errors (cm, ${}^{\circ}$) of standard descriptor-based localization techniques and descriptor-free methods, including the proposed DGC-GNN. The proposed method consistently outperforms GoMatch on all scenes by a significant margin. On the Cambridge-Landmarks dataset, the average median pose error of DGC-GNN is 54 cm / ${}^{\circ}$, while GoMatch leads to 173 cm / ${}^{\circ}$ average error. On 7Scenes, the average error of DGC-GNN is 15 cm / ${}^{\circ}$ and that of GoMatch is 22 cm / ${}^{\circ}$. DGC-GNN requires a similar amount of memory to other descriptor-free methods. Also, it inherits their privacy-preserving properties due to not requiring and storing visual descriptors.
The difference between descriptor-based (DB) and descriptor-free (DF) algorithms is clear. While descriptor-based ones lead to the best accuracy overall, they require excessive memory and descriptor maintenance and are susceptible to privacy attacks. Although the model compression method, HybridSC camposeco2019hybrid , shows effectiveness in storage saving, it achieves similar performance compared to DGC-GNN on the Cambridge dataset while still requiring descriptor maintenance. End-to-end methods (E2E) overcome these problems and achieve accurate results. However, their main limitation is that such approaches must be trained independently on each scene. The proposed DGC-GNN only needs to be trained once, making it more efficient and convenient to use as an off-the-shelf tool.
4 Conclusion
In conclusion, this paper introduces DGC-GNN, a novel graph-based pipeline for descriptor-free 2D-3D matching that effectively leverages geometric and color cues in a global-to-local manner. Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming a geometric embedding to guide the local point matching. By employing a cluster-based transformer, we enable efficient information passing within local clusters, ultimately leading to significant improvements in the number of correct matches and the accuracy of pose estimation. Compared to the state-of-the-art descriptor-free matcher GoMatch zhou2022geometry , the proposed DGC-GNN demonstrates a substantial improvement, doubling the accuracy on real-world and large-scale datasets. Furthermore, it results in significantly increased localization accuracy. These advancements contribute to reducing the gap between descriptor-based and descriptor-free methods while addressing the limitations of descriptor-based ones, such as memory footprint, maintenance costs, and susceptibility to privacy attacks. The source code and models will be made publicly available.
Limitations. The primary limitation of our proposed DGC-GNN method lies in its performance being inferior to traditional descriptor-based algorithms. The performance difference can be attributed to the insufficiency of unique 3D structures in the geometry, which hinders the ability of the algorithm to identify distinct matches in real-world scenarios. Although DGC-GNN demonstrates a notable improvement over existing descriptor-free approaches, there remains a performance gap to overcome in order to achieve results on par with or superior to those of descriptor-based methods.
Acknowledgements. This work was supported by the Academy of Finland (grant No. 327911, No. 353138), and by ETH Postdoc fellowship . We acknowledge the computational resources provided by the Aalto Science-IT project and CSC-IT Center for Science, Finland.
5 Supplementary Material
Appendix A Training and Evaluation Details
Dataset Generation. The training data generation process for MegaDepth li2018megadepth follows the methodology outlined in GoMatch zhou2022geometry . The undistorted SfM model reconstructions used in MegaDepth are provided by D2Net dusmanu2019d2 . For training, we sample up to 500 images from each scene. For each sampled image, we select the top-$k$ co-visible views that have at least 35% image overlap. This ensures that there are enough matches available for training. The overlapping score is computed by dividing the number of co-visible 3D points by the total number of points in the training image.
In the case of ScanNet dai2017scannet , a similar procedure is conducted. We also sample up to 500 images from each scene for the training set generation. The co-visible images are obtained using the co-visible scores provided by LOFTR sun2021loftr . We extract all the co-visible views of the training image with co-visible scores larger than 0.65, and then randomly sample the top-$k$ views for training. Since ScanNet is an RGB-D dataset without SfM models, we obtain the 3D points for each image by projecting the detected 2D keypoints with valid depth to the 3D coordinate. By doing this for each image, we reconstruct a sparse 3D point cloud based on the detected 2D keypoints. Note that the correspondence between different co-visible frames is not required in this case.
In total, for MegaDepth, we generate a training set consisting of 25,624 images from 99 scenes and a test set comprising 12,399 images covering 53 scenes. For ScanNet, we create a training set with 52,008 images from 105 scenes. The test set for ScanNet consists of 14,892 query images from 30 scenes. The data generation of 7Scenes shotton2013scene and Cambridge Landmarks kendall2015posenet follows the same procedure in  zhou2022geometry .
Inference. We consider a query with at least 10 keypoints as valid input. The 3D points from the top-$k$ retrieved database images are then applied to match against the queries with our proposed pipeline. We use the Sinkhorn algorithm sinkhorn1967concerning ; cuturi2013sinkhorn to optimize the extended cost matrix ¯$\bar{\mathcal{M}}\in\mathbb{R}^{N+1,M+1}$ in an iterative manner with up to 20 iterations to obtain the initial matches. The final matches are obtained by filtering the matches with matching confidence $\theta<0.5$ in the outlier rejection module. For the visual localization task, the camera poses are estimated by the P3P solver with RANSAC fischler1981random implemented in OpenCV bradski2000opencv and then refined by Levenberg-Marquardt more2006levenberg algorithm on the inliers matches, minimising the reprojection error.
Figure 5: 2D-3D Matching (shown by green lines) with the proposed DGC-GNN and GoMatch zhou2022geometry .
Appendix B Additional Results
Qualitative Results. More visualizations of inlier matches provided by DGC-GNN and GoMatch on MegaDepth are shown in Fig. 5. DGC-GNN consistently achieves more matches on multiple scenes, highlighting the effectiveness of the proposed method.
Methods Global C. Att. Color Ang. Cluster Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
@1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
GoMatch zhou2022geometry 15.67 / 22.43 / 28.01 0.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73
Variants G.Emb K-means 17.68 / 28.41 / 34.36 0.28 / 16.78 / 34.52 0.03 / 0.73 / 3.77
G.Label K-means 17.13 / 27.33 / 33.18 0.31 / 17.34 / 33.63 0.03 / 0.76 / 3.64
G.Emb K-means 18.10 / 30.64 / 37.07 0.24 / 14.48 / 34.30 0.03 / 0.63 / 3.51
G.Emb K-means 19.82 / 35.29 / 41.16 0.17 / 12.88 / 31.74 0.02 / 0.27 / 3.24
G.Emb Mean-shift 10.07 / 36.01 / 43.03 0.16 / 12.15 / 28.99 0.01 / 0.20 / 3.26
DGC-GNN G.Emb K-means 10.20 / 37.64 / 44.04 0.15 / 11.53 / 27.93 0.01 / 0.15 / 3.00
Table 4: Additional Ablation Results. AUC scores thresholded at 1, 5, and 10 pixels on $k=1$; rotation and translation error quantiles at 25, 50, 75% with the proposed components added one by one to the GoMatch pipeline.
(a) GoMatch
(b) +G. Emb, +Cluster Attention
(c) +G. Emb, +Cluster Attention, +Color
(d) DGC-GNN
Figure 6: Qualitative Matching Results of Different Architectures. We visualize the number of inlier matches after the PnP-RANSAC with different architectures (shown by green lines ).
Figure 7: Architectural Ablation Study. We plot the training loss and the number of matches estimated during the training for different architectures. Notably, we observe that the inclusion of color encoding in the training procedure leads to a significant improvement.
Additional Ablation Results. In addition to the ablation results presented in the main paper, we also provide ablation results for single-view matching with $k=1$ on MegaDepth li2018megadepth . Furthermore, we conduct two additional ablations to investigate the impact of different component selections. Firstly, we compare the effectiveness of the geometric global embedding (G. Emb.) used in the main paper with the global clustering label embedding (G. Label). Instead of encoding geometric cues, we encode the label of each global cluster and concatenate it to the local point feature. Then, we explore the selection of different clustering algorithms. We compare the performance of K-Means and Mean-Shift clustering algorithms in our pipeline.
The results are presented in Table 4. We observe similar conclusions for each component as in the main paper. The results obtained using the global label embedding (G. Label) with cluster attention (C.Att). show even worse performance compare to geometric embedding (G. Emb.) only, indicating the superiority of our clustering-based geometric embedding over the label embedding and highlighting the importance of incorporating geometric cues in the embedding process for effective point matching. Regarding the impact of different clustering algorithms, we only observe a minor difference in K-Means and Mean-Shift results, suggesting that our approach is robust to the choice of the clustering algorithm.
In addition to the numerical results, we also visualize the inlier matches (see Fig. 6), the training loss and the number of predicted matches during training (see Fig. 7) with different architectures to provide deep insights into the behaviour and performance of different architectures.
Appendix C Generalizability
Similar to zhou2022geometry , we discuss the generalizability of our DGC-GNN model on the visual localization task across different types of training and evaluation scenes. Specifically, we investigate the performance of our model when trained on MegaDepth li2018megadepth and ScanNet dai2017scannet , and evaluated on the 7Scenes dataset shotton2013scene . We also explore the impact of using different keypoint detectors, namely SIFT lowe2004distinctive and SuperPoint detone2018superpoint , during the evaluation.
The results of these experiments are summarized in Table 5, which provides an overview of the performance of our DGC-GNN model under different training and evaluation conditions. The results show that the DGC-GNN model trained on the ScanNet dataset, using the SIFT keypoint detector, and evaluated on the indoor 7Scenes dataset with the same detector, achieves the best overall performance. This demonstrates the effectiveness and robustness of our model in this specific training and evaluation scenario. However, we also note minor differences in performance when the model is trained on the MegaDepth dataset or evaluated using SuperPoint keypoints. Despite these variations, our findings suggest that the DGC-GNN model is scene-agnostic and can generalize well to different types of keypoint detectors. This is an encouraging result as showcases the versatility and generalization capability of the DGC-GNN approach, allowing it to be applied in various settings without significant degradation in performance.
Scenes Trained on MegaDepth li2018megadepth (SIFT) Trained on ScanNet dai2017scannet (SIFT)
GoMatch zhou2022geometry (SIFT) GoMatch zhou2022geometry (SP) DGC-GNN (SIFT) DGC-GNN (SP) DGC-GNN (SIFT) DGC-GNN (SP)
Chess 14 / 11.65 14 / 11.56 13 / 11.41 13 / 11.46 13 / 11.43 14 / 11.51
Fire 13 / 13.86 12 / 13.71 15 / 11.81 17 / 12.30 15 / 11.77 16 / 12.03
Heads 19 / 15.17 15 / 13.43 15 / 13.13 14 / 12.78 14 / 12.95 14 / 13.02
Office 11 / 12.48 17 / 11.76 17 / 11.66 17 / 11.66 16 / 11.61 17 / 11.66
Pumpkin 16 / 13.32 28 / 15.65 18 / 12.03 12 / 12.75 18 / 11.93 10 / 12.38
Redkitchen 13 / 12.84 14 / 13.03 18 / 12.14 10 / 12.36 18 / 12.09 19 / 12.28
Stairs 89 / 21.12 58 / 13.12 83 / 21.53 55 / 13.05 71 / 19.50 58 / 14.32
Table 5: Model Generalizability on Visual Localization Task. We report the translation and rotation median error (cm / ${}^{\circ}$) on 7Scenes dataset shotton2013scene . We evaluate the model trained on both MegaDepth li2018megadepth and ScanNet dai2017scannet datasets. The best performance is bold.
Appendix D Model Parameters and Timing
We discuss the model parameters and running time of DGC-GNN in this section. DGC-GNN incorporate global geometric embedding and local clustering attention, which has around 5.7 million trainable parameters and an estimated model size of 22.6 MB. The average inference time for each image pair over the Megadepth evaluation queries is 77.8ms, and the number is 70.9ms on ScanNet queries. The measurements are conducted on a 32GB NVIDIA Telsa V100 GPU.

### v2

@@
-DGC-GNN: Descriptor-free Geometric-Color Graph Neural Network for 2D-3D Matching
+DGC-GNN: Leveraging Geometry and Color Cues for Visual Descriptor-Free 2D-3D Matching
 Abstract
-Direct matching of 2D keypoints in an input image to a 3D point cloud of the scene without requiring visual descriptors has garnered increased interest due to its lower memory requirements, inherent privacy preservation, and reduced need for expensive 3D model maintenance compared to visual descriptor-based methods.
+Matching 2D keypoints in an image to a sparse 3D point cloud of the scene without requiring visual descriptors has garnered increased interest due to its low memory requirements, inherent privacy preservation, and reduced need for expensive 3D model maintenance compared to visual descriptor-based methods.
 However, existing algorithms often compromise on performance, resulting in a significant deterioration compared to their descriptor-based counterparts.
-In this paper, we introduce DGC-GNN, a novel algorithm that employs a global-to-local Graph Neural Network (GNN) that progressively exploits geometric and color cues to represent keypoints, thereby improving matching robustness.
-Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming the geometric embedding to guide the local point matching.
-We evaluate DGC-GNN on both indoor and outdoor datasets, demonstrating that it not only doubles the accuracy of the state-of-the-art descriptor-free algorithm but, also, substantially narrows the performance gap between descriptor-based and descriptor-free methods.
-The code and trained models will be made publicly available.
+In this paper, we introduce DGC-GNN, a novel algorithm that employs a global-to-local Graph Neural Network (GNN) that progressively exploits geometric and color cues to represent keypoints, thereby improving matching accuracy.
+Our procedure encodes both Euclidean and angular relations at a coarse level, forming the geometric embedding to guide the point matching.
+We evaluate DGC-GNN on both indoor and outdoor datasets, demonstrating that it not only doubles the accuracy of the state-of-the-art visual descriptor-free algorithm but also substantially narrows the performance gap between descriptor-based and descriptor-free methods.
+[^1] The code and trained models are available at: https://github.com/AaltoVision/DGC-GNN-release.
 1 Introduction
-Establishing 2D-3D matches plays a crucial role in various computer vision applications, including visual localization irschara2009structure ; sattler2012improving ; sattler2018benchmarking ; sarlin2019coarse ; wang2021continual ; sarlin2020superglue , 3D reconstruction snavely2006photo ; cui2015global ; schoenberger2016sfm ; lindenberger2021pixel , and Simultaneous Localization and Mapping (SLAM) eade2006scalable ; mur2015orb ; mur2017orb .
+Establishing 2D-3D matches plays a crucial role in various computer vision applications, including visual localization [20, 40, 42, 37, 53, 38], 3D reconstruction [48, 8, 43, 25], and Simultaneous Localization and Mapping (SLAM) [15, 31, 30].
 Traditional methods for establishing point-to-point matches involve extracting keypoints and descriptors from a query image, then matching the 2D and 3D descriptors using exhaustive search.
-To circumvent the computationally expensive matching process, some approaches irschara2009structure ; sarlin2019coarse narrow the search space by employing image retrieval methods nister2006scalable ; Arandjelovic16 first to identify the most similar images in the database, and then perform descriptor-based image matching lowe2004distinctive ; detone2018superpoint ; dusmanu2019d2 ; sarlin2020superglue ; shi2022clustergnn between the query and retrieved images.
+To circumvent the computationally expensive matching process, some approaches [20, 37] narrow the search space by employing image retrieval methods [33, 1] first to identify the most similar images in the database, and then perform descriptor-based image matching [27, 11, 14, 38, 45] between the query and retrieved images.
 The 2D-3D correspondences are subsequently established by connecting the 2D-2D image matches with the prebuilt 2D-3D correspondences in the database.
-Another approach sattler2011fast proposes an efficient vocabulary-based method to directly build 2D-to-3D matches by searching through all point descriptors.
+Another approach [39] is to build 2D-to-3D matches by searching through all point descriptors with an efficient vocabulary-based method.
 Sattler et al.
-sattler2012improving ; sattler2016efficient further explore the combination of both 2D-3D and 3D-2D search as an active correspondence search step for a faster and more efficient matching process.
-While descriptor-based algorithms achieve state-of-the-art accuracy, they store and maintain high-dimensional visual descriptors for each point in potentially large 3D point clouds, significantly increasing both memory footprint and maintenance costs.
-The stored model often requires orders of magnitude more storage than the point cloud and images alone zhou2022geometry .
-These methods are susceptible to privacy attacks dosovitskiy2016inverting ; dosovitskiy2016generating ; chelani2021privacy and necessitate computationally expensive model maintenance and descriptor update procedures zhou2022geometry when incorporating new descriptors or points into the model.
+[40, 41] further explore the combination of both 2D-3D and 3D-2D search as an active correspondence search step for a faster and more efficient matching process.
+Figure 1: 2D-3D matching (shown by green lines) with the proposed DGC-GNN and GoMatch [56].
+In this example, DGC-GNN obtains 78 correct matches with 0.02 meters camera translation and ${}^{\circ}$ rotation errors, while GoMatch finds only 17 inliers with a pose error of 0.37 meters and ${}^{\circ}$.
+While descriptor-based algorithms achieve state-of-the-art accuracy, they store and maintain high-dimensional visual descriptors for each point in potentially large 3D point clouds.
+The stored model often requires orders of magnitude more storage than the point cloud and images alone [56].
+These methods are susceptible to privacy attacks [13, 12, 7, 34] and necessitate computationally expensive model maintenance and descriptor update procedures [56] when incorporating new descriptors or points into the model.
 Several approaches have been proposed to address these limitations.
-Yang et al.Yang_2022_CVPR employ a learned point selection module to sample a subset of the point cloud for scene compression.
-Other methods li2020hierarchical ; brachmann2021dsacstar directly learn a function that maps 2D pixels to 3D coordinates without explicitly storing the 3D scene.
-Additionally, ng2022ninjadesc introduces an adversarial learning framework to develop content-concealing visual descriptors that prevent privacy leakage.
-Recently, researchers campbell2020solving ; liu2020learning have begun exploring deep learning techniques for cross-domain direct 2D-3D matching and pose estimation without visual descriptors, showcasing the potential of descriptor-free matching through differentiable geometric optimization.
-The recently proposed GoMatch zhou2022geometry represents significant progress in descriptor-free 2D-3D matching, achieving reasonable matching performance on a variety of real-world datasets li2018megadepth ; shotton2013scene ; kendall2015posenet .
+Yang et al.
+[54] employ learned point selection to sample a subset of the point cloud for scene compression.
+Other methods [23, 3] directly learn a function that maps 2D pixels to 3D coordinates without explicitly storing the 3D scene.
+Additionally, [32] introduces an adversarial learning framework to develop content-concealing descriptors that prevent privacy leakage.
+Recently, researchers [5, 26] have begun exploring deep learning techniques for cross-domain direct 2D-3D matching and pose estimation without visual descriptors, showcasing the potential of descriptor-free matching through differentiable geometric optimization.
+The recently proposed GoMatch [56] represents significant progress in descriptor-free 2D-3D keypoints matching, achieving reasonable matching performance on a variety of real-world datasets [24, 46, 22].
 GoMatch first identifies keypoints in the query image, which, along with the 3D points from the model, are converted to bearing vectors in the camera coordinate system.
-The algorithm employs the attention mechanism sarlin2020superglue ; vaswani2017attention to effectively establish reliable 2D-3D correspondences.
-While GoMatch attains reasonable accuracy, its performance still significantly lags behind its descriptor-based counterparts sarlin2020superglue ; sarlin2019coarse ; sattler2016efficient .
-Additionally, our experiments reveal that its reliance on geometric cues from the points and their local neighbors renders GoMatch incapable of distinguishing geometrically indistinct structures.
-Figure 1: 2D-3D matching (shown by green lines) with the proposed DGC-GNN and GoMatch zhou2022geometry .
-In this example, DGC-GNN obtains 78 correct matches with 0.02 meters camera translation and ${}^{\circ}$ rotation errors, while GoMatch finds only 17 inliers with a pose error of 0.37 meters and ${}^{\circ}$.
+The algorithm employs an attention mechanism [38, 52] to establish reliable 2D-3D correspondences effectively.
+While GoMatch attains reasonable accuracy, its performance still significantly lags behind its descriptor-based counterparts [38, 37, 41].
+Additionally, it relies on geometric cues only from the points and their local neighbors, rendering it incapable of distinguishing geometrically similar structures.
 These observations lead us to two critical questions: (1) Is geometry the only information we can utilize?
@@
 Building upon these observations and the groundwork set by GoMatch, we propose a novel graph-based pipeline, named DGC-GNN, which leverages geometric and color cues in a global-to-local manner for descriptor-free 2D-3D matching.
-Our pipeline encodes both position and RGB information for each point and extracts a global distance-angular embedding to guide local point matching.
-Taking inspiration from shi2022clustergnn , we employ a cluster-based transformer to constrain information flow within local clusters.
-We observe, from real-world datasets, that the proposed DGC-GNN leads to substantial improvements in the number of correct matches and the accuracy of pose estimation.
+DGC-GNN encodes position and RGB information for each point and extracts a global distance-angular embedding to guide local point matching.
+Taking inspiration from [45], we employ a cluster-based transformer to constrain information flow within local clusters.
+We observe, from real-world datasets, that DGC-GNN leads to substantial improvements in the number of correct matches and the accuracy of pose estimation.
 Notably, it doubles the accuracy of GoMatch, thereby reducing the gap between descriptor-based and descriptor-free methods.
 In summary, our paper makes the following contributions:
-  • We introduce a descriptor-free global-to-local GNN for direct 2D-3D point matching.
+  • We introduce a visual descriptor-free global-to-local GNN for direct 2D-3D keypoint matching.
 The network leverages multiple cues and incorporates a progressive clustering module to represent the keypoints.
-This pipeline enhances the robustness and accuracy of 2D-3D matching while requiring low memory, being privacy-preserving and free from 3D model maintenance.
-  • We demonstrate that color information is crucial for 2D-3D matching. By incorporating RGB encoding into our network, we observe significant improvements in performance.
-  • Extensive experiments on multiple real-world datasets show that the proposed network outperforms previous methods by a large margin on both matching and visual localization tasks.
-2 Descriptor-Free 2D-3D Matching
+This pipeline enhances the accuracy of sparse 2D-3D matching while requiring low memory, being privacy-preserving, and low cost from 3D model maintenance.
+  • We demonstrate that color information for each point is crucial for 2D-3D matching. By incorporating RGB encoding into our network, we observe significant performance improvements.
+  • Extensive experiments on real-world datasets show that DGC-GNN outperforms previous methods by a large margin on both matching and visual localization tasks.
+Figure 2: Pipeline overview.
+For keypoints from the 2D image and 3D points from the point cloud, the proposed DGC-GNN (1) considers the bearing vectors and the color at each bearing vector as input.
+(2) It extracts the point-wise position and color features with two separate encoders and mixes the features as $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$.
+(3) The bearing vectors are clustered into $K$ groups, and geometric graphs are built upon the clusters to extract the global-level geometric embeddings $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$.
+(4) We then concatenate $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ with $\mathbf{f_{p}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$ with $\mathbf{f_{p}}$, and build a local graph at each point as self-attention.
+A cluster-based attention module is adopted to enhance the local features by forcing the message to pass only with the most related features.
+A differentiable layer matches and optimizes the improved features to obtain score matrix $\mathcal{S}$.
+Finally, an outlier rejection network is applied to prune the matches with low confidence, leading to the final 2D-3D correspondences $\mathcal{M}_{final}$.
+2 Visual Descriptor-Free 2D-3D Matching
 2.1 Problem Formulation and Notation
-Considering keypoints $\mathbf{P}=\{\mathbf{p}_{n}\in\mathbb{R}^{2}\;|\;n=1,...,N\}$ from a query image $I$ and a database 3D point cloud $\mathbf{Q}=\{\mathbf{q}_{m}\in\mathbb{R}^{3}\;|\;m=1,...,M\}$, where, optionally, each 3D point is associated with a visual descriptor $\mathbf{d}\in\mathbb{R}^{D}$.
-The task of 2D-3D point matching is to find a set $\mathcal{M}_{\mathbf{p,q}}$ of corresponding keypoints such that
+Given keypoints $\mathbf{P}=\{\mathbf{p}_{n}\in\mathbb{R}^{2}\;|\;n=1,...,N\}$ from query image $I$ and database 3D point cloud $\mathbf{Q}=\{\mathbf{q}_{m}\in\mathbb{R}^{3}\;|\;m=1,...,M\}$, where, optionally, each 3D point is associated with a visual descriptor $\mathbf{d}\in\mathbb{R}^{D}$.
+The task is to find a set $\mathcal{M}_{\mathbf{p,q}}$ of corresponding keypoints such that
 $\mathcal{M}_{\mathbf{p,q}}=\{(n,m)\;|\;||\pi(\mathbf{q}_{m},\mathbf{R,t,K})-\mathbf{p}_{n}||_{2}\leq\epsilon\},$ (1)
@@
 Bearing Vector.
-Similar to zhou2022geometry , we adopt bearing vectors as keypoint representation for both the 2D and 3D points to alleviate their cross-domain nature and represent them in the same space.
+Similar to [56], we adopt bearing vectors as keypoint representation for both the 2D and 3D points to alleviate their cross-domain nature and represent them in the same space.
 The bearing vector is the direction from the camera center to a 3D point in the camera coordinate system.
-Given an image, a 2D pixel $\mathbf{p}_{n}$ is uplifted to bearing vector as $[\mathbf{b}_{\mathbf{p},n},1]^{\text{T}}=\mathbf{K^{-1}}[\mathbf{p}_{n},1]^{\text{T}},\mathbf{b}_{\mathbf{p},,}\in\mathbb{R}^{2}$, where $\mathbf{K}$ is the intrinsic camera matrix.
-Given a 3D point $\mathbf{q}_{m}$, the corresponding bearing vector is $[\mathbf{b}_{\mathbf{p},m},1]^{\text{T}}=\frac{\mathbf{Rq}_{m}+\mathbf{t}}{[\mathbf{Rq}_{m}+\mathbf{t}]_{z}}$, where $\mathbf{R}$ is the camera rotation and $\mathbf{t}$ is its translation in the world coordinate system and subscript $z$ denotes the third component of the 3D vector.
+Given an image, a 2D pixel $\mathbf{p}_{n}$ is uplifted to bearing vector as $[\mathbf{b}_{\mathbf{p},n},1]^{\text{T}}=\mathbf{K^{-1}}[\mathbf{p}_{n},1]^{\text{T}},\mathbf{b}_{\mathbf{p},n}\in\mathbb{R}^{2}$, where $\mathbf{K}$ is the intrinsic camera matrix.
+Given a 3D point $\mathbf{q}_{m}$, the corresponding bearing vector is
+$[\mathbf{b}_{\mathbf{q},m},1]^{\text{T}}=\frac{\mathbf{Rq}_{m}+\mathbf{t}}{[\mathbf{Rq}_{m}+\mathbf{t}]_{z}},$ (2)
+where $\mathbf{R}$ is the camera rotation and $\mathbf{t}$ is its translation in the world coordinate system and subscript $z$ denotes the third component of the 3D vector.
 2.2 Network Architecture
-Figure 2: Pipeline overview.
-For keypoints from the 2D image and 3D points from the point cloud, the proposed DGC-GNN (1) considers the bearing vectors and the color at each bearing vector as input.
-(2) It extracts the point-wise position and color features with two separate encoders and mixes the features as $\mathbf{f_{p}}$ and $\mathbf{f_{p}}$.
-(3) The bearing vectors are clustered into $K$ groups, and geometric graphs are built upon the clusters to extract the global-level geometric embeddings $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$.
-(4) We then concatenate $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ with $\mathbf{f_{p}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$ with $\mathbf{f_{p}}$, and build a local graph at each point as self-attention.
-A cluster-based attention module is adopted to enhance the local features by forcing the message passing only with the most related features.
-A differentiable layer matches and optimizes the improved features to obtain score matrix $\mathcal{S}$.
-Finally, an outlier filtering network is applied to prune the matches with low confidence, leading to the final 2D-3D correspondences $\mathcal{M}_{final}$.
-The proposed DGC-GNN applies a hierarchical mechanism to leverage both color and geometric cues in a global-to-local fashion.
+The proposed DGC-GNN applies a hierarchical mechanism to leverage color and geometric cues in a global-to-local fashion.
 The overall pipeline is illustrated in Fig.2.
@@
 To extract points-wise features from both the 2D keypoint set $\mathbf{P}$ and the 3D point cloud $\mathbf{Q}$, we consider the inputs as bearing vectors equipped with color information: $\mathcal{P}=\{\mathbf{b_{p},c_{p}}\}$ and $\mathcal{Q}=\{\mathbf{b_{q},c_{q}}\}$.
-Two ResNet-style point encoders he2016deep ; campbell2020solving , denoted as $\mathcal{F}_{b}$ and $\mathcal{F}_{c}$, are applied to extract position and color embeddings separately.
+Two ResNet-style point encoders [18, 5], denoted as $\mathcal{F}_{b}$ and $\mathcal{F}_{c}$, are applied to extract position and color embeddings separately.
 We then obtain the local point features, $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$, as follows:
 $\mathbf{f_{p}}=\mathcal{F}_{b}(\mathbf{b_{p}})+\mathcal{F}_{c}(\mathbf{c_{p}}),\ \ \mathbf{f_{q}}=\mathcal{F}_{b}(\mathbf{b_{q}})+\mathcal{F}_{c}(\mathbf{c_{q}}).$ (3)
-The resulting point-wise features $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$ are vectors with dimensions $\mathbb{R}^{N\times d}$ and $\mathbb{R}^{M\times d}$ respectively, where $N$ and $M$ represent the number of keypoints in $\mathbf{P}$ and $\mathbf{Q}$, and $d$ denotes the dimensionality of the encoded features, e.g., $d=128$ for SIFT features lowe2004distinctive .
+The resulting point-wise features $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$ are vectors with dimensions $\mathbb{R}^{N\times d}$ and $\mathbb{R}^{M\times d}$ respectively, where $N$ and $M$ represent the number of keypoints in $\mathbf{P}$ and $\mathbf{Q}$, and $d$ denotes the dimensionality of the encoded features, e.g., $d=128$
 2.2.2 Global Geometric Guidance
-Global context guidance has demonstrated its effectiveness in various computer vision tasks sun2021loftr ; li2020hierarchical ; yu2021cofinet ; qin2022geometric .
+Global context guidance has demonstrated its effectiveness in various computer vision tasks [49, 23, 55, 36].
 Global context helps to differentiate local descriptors from similar structures or patches, thereby reducing ambiguity.
-However, most existing methods sun2021loftr ; qin2022geometric consider the outputs from different encoding layers as global and local features.
+However, most existing methods [49, 36] consider the outputs from different encoding layers as global and local features.
 This approach is not suitable for our scenario, as our input is sparse points.
@@
 As shown in Fig.
-3.(a) and (c), the input bearing vectors, both in the image and in the point cloud, are first clustered into $K$ groups.
-The groups represent distinct clusters, each associated with a cluster center as the global position, denoted by $\mathbf{\hat{b}}_{\mathbf{p},k}\in\mathbb{R}^{2},k=1,...K$.
-The corresponding global embedding is obtained by computing the average of the point embeddings within the cluster, thus $\mathbf{\hat{f}}_{\mathbf{p},k}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{f}_{\mathbf{p},k},\ \mathbf{\hat{f}}_{\mathbf{p},k}\in\mathbb{R}^{d}$.
+3.(a) and (c), the input bearing vectors, both in the image and in the point cloud, are first clustered into $X$ groups.
+The groups represent distinct clusters, each associated with a cluster center as the global position, denoted by $\mathbf{\hat{b}}_{\mathbf{p},x}\in\mathbb{R}^{2},x=1,...X$.
+The corresponding global embedding is obtained as the average of the point embeddings within a cluster as $\mathbf{\hat{f}}_{\mathbf{p},x}=\frac{1}{P^{\prime}}\sum_{p^{\prime}=1}^{P^{\prime}}\mathbf{f}_{\mathbf{p},p^{\prime}}$, where $\mathbf{\hat{f}}_{\mathbf{p},x}\in\mathbb{R}^{d}$, $P^{\prime}$ is the number of points in the $x$.
 The same is conducted on the 3D points to obtain $\mathbf{\hat{b}}_{\mathbf{q}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}$.
@@
 Global Geometric Graph.
-To aggregate and extract the geometric relations among the clusters, we propose a novel graph neural network that encodes both distance and angular cues; the basic GNN structure is built upon huang2021predator ; zhou2022geometry .
-In the following, we describe the graph building for the 2D global points set $\mathcal{\hat{P}}=\{\mathbf{\hat{b}_{p}},\mathbf{\hat{f}_{p}}\}$ and the same goes for $\mathcal{\hat{Q}}=\{\mathbf{\hat{b}_{q}},\mathbf{\hat{f}_{q}}\}$.
-Each cluster center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ is connected to its $k$-NN neighbours ($k\leq K$) in the coordinate space, and $\xi_{\mathbf{p},(x,y)}$ is the edge between center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$.
+To aggregate and extract the geometric relations among the clusters, we propose a novel graph neural network that encodes both distance and angular cues; the basic GNN structure is built upon [19, 56].
+In the following, we describe the graph construction for the 2D global points set $\mathcal{\hat{P}}=\{\mathbf{\hat{b}_{p}},\mathbf{\hat{f}_{p}}\}$ and the same goes for $\mathcal{\hat{Q}}=\{\mathbf{\hat{b}_{q}},\mathbf{\hat{f}_{q}}\}$.
+Each cluster center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ is connected to its $k$-NN neighbours ($k\leq K$) in the coordinate space, and $\xi_{\mathbf{p},(x,y)}$ is the edge between center points $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$.
 We update the feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ using the following equation:
-${}^{(t+1)}\mathbf{\hat{f}}_{\mathbf{p},x}=\max_{\xi_{\mathbf{p},(x,y)}}\mathcal{H}_{g}(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}-^{(t)}\mathbf{\hat{f}}_{\mathbf{p},y})),$ (3)
-where the $\oplus$ denotes concatenation and $\mathcal{H}_{g}(*)$ is the linear projection with instance normalization ulyanov2016instance and a LeakyReLU function maas2013rectifier .
+${}^{(t+1)}\mathbf{\hat{f}}_{\mathbf{p},x}=\max_{\xi_{\mathbf{p},(x,y)}}\mathcal{H}_{g1}(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}-^{(t)}\mathbf{\hat{f}}_{\mathbf{p},y})),$ (4)
+where the $\oplus$ denotes concatenation and $\mathcal{H}_{g1}(*)$ is the linear projection with instance normalization [51] and a LeakyReLU function [28].
+The $max$ operator applies to the $k$-NN neighbors.
 The global feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ is updated twice, and calculated as
-$\mathbf{\hat{f}}_{\mathbf{p},x}^{g}=\mathcal{H}_{g}(^{(0)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(1)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(2)}\mathbf{\hat{f}}_{\mathbf{p},x}).$ (4)
-Besides the distance embedding, inspired by qin2022geometric , we also adopt the angular embedding to obtain rotation-invariant geometric cues for the global point representation.
+$\mathbf{\hat{f}}_{\mathbf{p},x}^{g}=\mathcal{H}_{g2}(^{(0)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(1)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(2)}\mathbf{\hat{f}}_{\mathbf{p},x}).$ (5)
+$\mathcal{H}_{g2}$ has a similar structure to $\mathcal{H}_{g1}$, but without shared weights.
+Besides the distance embedding, inspired by [36], we also adopt the angular embedding to obtain rotation-invariant geometric cues for the global representation.
 To do so, we define the embedding on cluster triplets as shown in Fig.
-3.(b).
-Given bearing vector $\mathbf{\hat{b}}_{\mathbf{p},z}$ and two of its neighbors $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$, the angular embedding of $\mathbf{\hat{b}}_{\mathbf{p},z}$ is defined as follows:
+3.
+(b).
+Given bearing vector $\mathbf{\hat{b}}_{\mathbf{p},x}$ and two of its neighbors $\mathbf{\hat{b}}_{\mathbf{p},y}$ and $\mathbf{\hat{b}}_{\mathbf{p},z}$, the angular embedding of <$\mathbf{\hat{b}}_{\mathbf{p},x}$, $\mathbf{\hat{b}}_{\mathbf{p},y}$> w.r.t.
+$\mathbf{\hat{b}}_{\mathbf{p},z}$ is defined as follows:
 $\mathbf{A}_{x,y}^{z}=\operatorname{sine}(\angle(\mathbf{\hat{b}}_{\mathbf{p},z}-\mathbf{\hat{b}}_{\mathbf{p},x},\ \ \ \mathbf{\hat{b}}_{\mathbf{p},y}-\mathbf{\hat{b}}_{\mathbf{p},x})/\sigma_{a}),$ (6)
-where $\operatorname{sine}(\cdot)$ is a sinusoidal function and $\sigma_{a}$ is a controller constant.
-We update the global geometric embedding $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ as an angular-aware attention mechanism as follows:
+where $\operatorname{sine}(\cdot)$ is a sinusoidal function and $\sigma_{a}$ is a controller constant, all $k$ neighbours are considered to obtain the angular embedding $\mathbf{A}_{\mathbf{p}}$.
+We update the global geometric embedding $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ as an angular-aware attention mechanism:
 $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}=\operatorname{norm}(\mathbf{\hat{f}}_{\mathbf{p}}^{g}+\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}}));\ \ \ \ \mathbf{\hat{f}}_{\mathbf{p}}^{gg}\in\mathbb{R}^{K\times d},$ (7)
 where
-$\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}})=(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{V}}).\frac{(\mathbf{A}_{\mathbf{p}}\mathbf{W^{A}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})^{\text{T}}+(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{K}})^{\text{T}}}{\sqrt{dim}}.$ (7)
-$\mathbf{W^{A}},\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{d\times d}$ are the projection matrices of each item.
-The local point representation with global geometric embedding is as $\mathcal{\widetilde{P}}=\{\mathbf{b}_{\mathbf{p}},\mathbf{\widetilde{f}}_{\mathbf{p}}\}$ and $\mathbf{\widetilde{f}}_{\mathbf{p}}=\mathbf{f}_{\mathbf{p}}\oplus\mathbf{\hat{f}}_{\mathbf{p}}^{gg},\mathbf{\widetilde{f}}_{\mathbf{p}}\in\mathbb{R}^{N\times 2d}$.
-The operation for the bearing vectors obtained from point cloud $\mathcal{\hat{Q}}$ is the same and we get $\mathcal{\widetilde{Q}}=\{\mathbf{b}_{\mathbf{q}},\mathbf{\widetilde{f}}_{\mathbf{q}}\}$.
+$\begin{split}\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}})=\\ (\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{V}}).\frac{(\mathbf{A}_{\mathbf{p}}\mathbf{W^{A}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})^{\text{T}}+(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{K}})^{\text{T}}}{\sqrt{dim}}.\end{split}$
+$\mathbf{W^{A}},\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{d\times d}$ are the projection matrices of each item and LayerNorm [2] is applied to Eq.
+7.
+Each local point feature is associated with its corresponding global embedding by $\mathbf{\widetilde{f}}_{\mathbf{p}}=\mathbf{f}_{\mathbf{p}}\oplus\mathbf{\hat{f}}_{\mathbf{p}}^{gg},\mathbf{\widetilde{f}}_{\mathbf{p}}\in\mathbb{R}^{N\times 2d}$ to obtain $\mathcal{\widetilde{P}}=\{\mathbf{b}_{\mathbf{p}},\mathbf{\widetilde{f}}_{\mathbf{p}}\}$.
+The same procedure obtains local and global embedding $\mathcal{\widetilde{Q}}=\{\mathbf{b}_{\mathbf{q}},\mathbf{\widetilde{f}}_{\mathbf{q}}\}$ for the point cloud $\mathcal{\hat{Q}}$.
 2.2.3 Cluster-based Local Matching
 After extracting the global geometric embedding, we implement a cluster-based matching module to obtain the initial intra-domain 2D-3D matches.
-This cluster-based GNN  shi2022clustergnn has been shown to be more computationally efficient than its complete-graph counterpart sarlin2020superglue .
-The network considers the local point features from both $\mathcal{\widetilde{P}}$ and $\mathcal{\widetilde{Q}}$ a complete set, then classifies the local feature with strong correlations into the same group and restricts the message passing within each group.
-In addition to its low computational complexity, we found that the cluster GNN can effectively utilize our global-to-local geometric cues, as the clustering operation inherits the property of global graph clustering and forces it to distinguish ambiguous local features even with similar global embedding.
-GoMatch-style Initialization.
-Our local graph initialization follows the pipeline in GoMatch zhou2022geometry and conducts it as geometric self-attention.
+This cluster-based GNN  [45] has been shown to be more computationally efficient than its complete-graph counterpart [38].
+The network considers the local point features from both $\mathcal{\widetilde{P}}$ and $\mathcal{\widetilde{Q}}$ a complete set, then clusters the feature with strong correlations into the same group and restricts the message passing within each group.
+In addition to its low computational complexity, we found that cluster GNN can effectively utilize global-to-local geometric cues, as the clustering operation inherits the property of global graph clustering and forces it to distinguish ambiguous local features even with similar global embedding.
+Methods ScanNet [10] MegaDepth [24]
+Reproj. AUC (%) Rotation (${}^{\circ}$) Translation (m) P (%) ($\uparrow$) Reproj. AUC (%) Rotation (${}^{\circ}$) Translation P (%) ($\uparrow$)
+@1 / 5 / 10px ($\uparrow$) Quantile @25 / 50 / 75% ()($\downarrow$)( ↓ ) @1 / 5 / 10px ($\uparrow$) Quantile @25 / 50 / 75% ()($\downarrow$)( ↓ )
+k=1 Oracle 29.13 / 39.83 / 41.34 110.10 / 110.19 / 110.40 0.01 / 0.01 / 0.03 - 34.59 / 85.02 / 92.02 10.04 / 10.06 / 10.12 0.00 / 0.01 / 0.01 -
+BPnPNet [5] 10.00 / 10.00 / 10.02 199.17 / 128.90 / 154.68 4.35 / 6.82 / 9.86 13.60 10.22 / 10.63 / 10.89 16.13 / 32.01 / 61.58 1.67 / 3.17 / 5.44 12.95
+GoMatch [56] 11.18 / 11.23 / 18.01 112.69 / 112.78 / 136.50 0.19 / 0.91 / 2.63 13.18 15.67 / 22.43 / 28.01 10.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73 14.94
+DGC-GNN 12.73 / 21.88 / 32.23 110.94 / 113.17 / 120.14 0.06 / 0.23 / 1.40 14.86 10.20 / 37.64 / 44.04 10.15 / 11.53 / 27.93 0.01 / 0.15 / 3.00 19.00
+k=10 BPnPNet [5] 10.00 / 10.00 / 10.03 104.68 / 135.94 / 160.54 4.67 / 7.30 / 10.92 10.84 10.36 / 10.72 / 10.97 16.63 / 34.69 / 67.77 1.64 / 3.30 / 5.97 10.74
+GoMatch [56] 10.91 / 18.98 / 31.12 111.18 / 114.94 / 128.97 10.08 / 0.35 / 2.08 14.25 18.90 / 35.67 / 44.99 10.18 / 11.29 / 16.65 0.02 / 0.12 / 1.92 18.76
+DGC-GNN 11.76 / 31.74 / 48.11 110.67 / 111.49 / 117.62 10.04 / 0.11 / 0.53 16.42 15.30 / 51.70 / 60.01 10.07 / 10.26 / 15.41 0.01 / 0.02 / 0.57 13.36
+Table 1: 2D-3D Matching.
+We present AUC scores for reprojection errors threshold at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, and 75%; and matching precision.
+Parameter $k$ is the number of images retrieved from the database to narrow down the search space.
+The best results are bold.
+DGC-GNN nearly doubles the AUC scores of GoMatch and reduces the pose errors to their ${\approx}33\%$.
+Initialization.
+As an initialization procedure for the cluster attention module, we run the general self and cross-attention modules proposed in GoMatch [56].
 For each local point $\mathbf{b}_{\mathbf{p},n}$, we construct a local graph according to its $k^{\prime}$ nearest neighbours in the Euclidean space and update the associated feature $\mathbf{\widetilde{f}}_{\mathbf{p},n}\in\mathbb{R}^{2d}$ by Eq.
-4.
+5.
 Note that we ignore the angular embedding at this stage due to the unaffordable memory requirements with space complexity $\mathcal{O}(Nk^{\prime 2})$, where $N$ is the number of local points.
-We then adopt linear attention katharopoulos2020transformers ; sun2021loftr as a cross-attention mechanism, which allows each keypoint in one modality to interact with all keypoints from another modality.
+We then use linear attention [21, 49] as a cross-attention mechanism, which allows each point in one modality to interact with all points from another modality.
 This not only facilitates inter-modality in the feature matching but also reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
@@
 After the graph initialization, the features $\mathbf{\widetilde{f}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}}_{\mathbf{q}}$ coming from the image and the point cloud respectively, are concatenated and processed in a two-level hierarchical clustering attention module.
-The hierarchical structure is effective in reducing erroneous grouping.
-At the first level, we directly cluster the feature vectors into $I$ coarse groups.
+The hierarchical structure is effective in suppressing erroneous groupings.
+At the first level, we cluster the feature vectors into $I$ coarse groups.
 In the second level, each coarse group is divided into several small groups.
@@
 […]
+$\mathcal{M}_{init}=\{(\widetilde{n},\widetilde{m})\ |\ \forall(\widetilde{n},\widetilde{m})\in\text{MNN}(\mathcal{S})\},$ (8)
+where MNN is the mutual nearest neighbors operator. Set $\mathcal{M}_{init}$ provides initial 2D-3D matches that we further filter in Sec. 2.3 to keep the accurate correspondences only.
 2.3 Outlier Rejection
-After obtaining the initial matches, an outlier pruning approach is adopted to remove the incorrect ones.
-We apply a classification network, as applied in yi2018learning ; zhou2022geometry , whose input is the concatenated 2D and 3D keypoint features $\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}=\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p},\widetilde{n}}\oplus\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q},\widetilde{m}}$ and outputs the matching confidence of each matched pair.
+After obtaining the initial matches, outlier pruning runs to remove the incorrect ones.
+We apply the same outlier rejection network as in GoMatch [56], whose input is the concatenated 2D and 3D keypoint features $\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}=\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p},\widetilde{n}}\oplus\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q},\widetilde{m}}$ and outputs the matching confidence of each matched pair.
 The final predicted matches are obtained as follows:
-$\mathcal{M}_{final}=\{(\widetilde{n}^{\prime},\widetilde{m}^{\prime})\ |\ \forall\ \operatorname{classifier}(\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}\;|\;(\widetilde{n},\widetilde{m})\in\mathcal{M}_{init})\geq\theta\},$ (8)
+$\mathcal{M}_{final}=\{(\widetilde{n}^{\prime},\widetilde{m}^{\prime})\ |\forall\ \text{cls}(\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}\;|\;(\widetilde{n},\widetilde{m})\in\mathcal{M}_{init})\geq\theta\},$ (9)
 where $\theta$ is the matching confidence threshold.
@@
 The matching loss $\mathcal{L}_{ot}$ minimizes the negative log-likelihood of the matching score $\mathcal{\bar{S}}$.
-$\mathcal{L}_{ot}=-\frac{1}{|\mathcal{M}_{gt}|+|\mathcal{I}|+|\mathcal{J}|}(\sum\limits_{(n,m)\in\mathcal{M}_{gt}}\log\mathcal{\bar{S}}_{n,m}+\sum\limits_{n\in\mathcal{I}}\log\mathcal{\bar{S}}_{i,m+1}+\sum\limits_{m\in\mathcal{J}}\log\mathcal{\bar{S}}_{N+1,j}).$ (9)
+$\begin{split}\mathcal{L}_{ot}=-\frac{1}{|\mathcal{M}_{gt}|+|\mathcal{I}|+|\mathcal{J}|}(\sum\limits_{(n,m)\in\mathcal{M}_{gt}}\log\mathcal{\bar{S}}_{n,m}+\\ \sum\limits_{i\in\mathcal{I}}\log\mathcal{\bar{S}}_{i,m+1}+\sum\limits_{j\in\mathcal{J}}\log\mathcal{\bar{S}}_{N+1,j}).\end{split}$ (10)
 The classification loss is defined as
 $\mathcal{L}_{or}=-\frac{1}{|\mathcal{M}_{init}|}\sum_{i=1}^{|\mathcal{M}_{init}|}w_{i}(y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})),$ (11)
-where $w_{i}$ is the balance weight of positive and negative samples, $y_{i}$ is the gt matching label for the $i$-th correspondences, $p_{i}$ is the predicted probability of a true match for the $i$-th correspondences.
-The total loss is the sum of the two terms, thus, $\mathcal{L}=\mathcal{L}_{ot}+\mathcal{L}_{or}$.
+where $w_{i}$ is the weight balancing the positive and negative samples, $y_{i}$ is the ground truth matching label for the $i$-th correspondences, $p_{i}$ is the predicted probability of a true match for the $i$-th correspondences.
+The total loss is the sum of the two terms as $\mathcal{L}=\mathcal{L}_{ot}+\mathcal{L}_{or}$.
 3 Experiments
-In this section, we first describe the experimental setup, employed datasets, and evaluation protocols.
-Next, we provide detailed comparisons with the baselines on both matching and visual localization tasks.
-Finally, we conduct ablation studies on each designed component.
 Training.
-We train the indoor model of DGC-GNN on the ScanNet dai2017scannet dataset and the outdoor model on the MegaDepth li2018megadepth dataset.
-We extract up to 1024 keypoints for each training image by the SIFT detector lowe2004distinctive .
-Similarly as in GoMatch, we first select a subset of the point cloud by applying image retrieval approaches Arandjelovic16 ; torii201524 to obtain potential images observing the same part of the scene as the input one.
-During the training, we randomly sample the retrieval pairs with a visual overlap of more than 35% on MegaDepth and 65% on ScanNet to ensure enough matches on each pair.
-For the global geometric embedding, we cluster the 2D/3D bearing vectors into $K=10$ groups, and each cluster center point is connected to its $k=4$ nearest neighbors to build the global graph.
+We train the indoor model of DGC-GNN on the ScanNet [10] dataset and the outdoor model on the MegaDepth [24] dataset.
+We extract up to 1024 keypoints for each training image by the SIFT detector [27].
+Similarly as in GoMatch, we first select a subset of the point cloud by applying image retrieval approaches [1, 50] to obtain potential images observing the same part of the scene as the input one.
+We randomly sample the retrieval pairs with a visual overlap of more than 35% on MegaDepth and 65% on ScanNet to ensure enough matches on each pair.
+For the global geometric embedding, we cluster the 2D/3D bearing vectors into $X=10$ groups, and each cluster center is connected to its $k=4$ nearest neighbors to build the global graph.
 For the local point graph, we connect each point with its 10 nearest neighbors and the cluster-based attentions are performed twice to force the intra-cluster information exchange.
-The model is optimized using the Adam with a fixed learning rate of 1e-3.
-We train DGC-GNN with one 32GB Telsa V100 GPU.
-The convergence of the model typically requires 25 epochs with around 20 hours for indoor scenes and 35 epochs (approx.
-35 hours) for outdoor scenes.
+We use Adam optimizer with a learning rate of 1e-3. We train DGC-GNN with one 32GB Telsa V100 GPU. The convergence of the model typically requires 50 epochs.
 Datasets.
-We use ScanNet and MegaDepth for the model training and 2D-3D matching task evaluation.
-As a downstream application, we perform visual localization on the 7Scenes shotton2013scene and Camdridge-Landmarks kendall2015posenet datasets.
+We use ScanNet and MegaDepth for training and 2D-3D matching task evaluation.
+As a downstream application, we perform visual localization on the 7Scenes [46] and Cambridge Landmarks [22] datasets.
 MegaDepth is a popular outdoor dataset with 196 scenes captured around the world.
-The sparse 3D reconstructions are provided by the COLMAP schoenberger2016sfm structure-from-motion software.
-Following zhou2022geometry , we train our outdoor model on 99 scenes and evaluate it with 53 scenes.
-ScanNet is a large-scale RGB-D indoor dataset that consists of 1613 scans with over 2.5 million images.
-We randomly selected 105 scenes for the training and 30 scenes for the evaluation.
-Cambridge-Landmarks is a middle-scale outdoor dataset consisting of 6 individual scenes.
-The ground truth camera poses are provided by a structure-from-motion algorithm.
-We follow kendall2015posenet ; zhou2022geometry to evaluate our method on four of the scenes.
+The sparse 3D reconstructions are provided by the COLMAP [43] structure-from-motion software.
+Following [56], we train our outdoor model on 99 scenes and evaluate it with 53 scenes.
+ScanNet is a large-scale RGB-D indoor dataset comprising 1613 scans with over 2.5 million images.
+We randomly selected 105 scenes for the training and 30 for the evaluation.
+Cambridge Landmarks is a middle-scale outdoor dataset consisting of 6 individual scenes.
+A structure-from-motion algorithm provides the ground truth camera poses.
+We follow [22, 56] to evaluate our method on four scenes.
 7Scenes is a small indoor dataset with RGB-D images and camera poses provided by the depth SLAM system.
@@
 Evaluation Protocol.
-For matching on ScanNet and MegaDepth, we follow zhou2022geometry and report the AUC score calculated from the reprojection errors.
-To calculate the reprojection errors for the 2D-3D matches in $\mathcal{M}_{final}$, we project the 3D points to the image plane using the ground truth and estimated camera poses.
-Then we calculate the $L_{2}$ distance between the ground truth and estimated reprojected 2D points.
-We use multiple thresholds, 1, 5, and 10 pixels, to evaluate the algorithms.
+For matching on ScanNet and MegaDepth, we follow [56] and report the AUC score calculated from the reprojection errors.
+To calculate the errors for the 2D-3D matches in $\mathcal{M}_{final}$, we project the 3D points to the image plane using the ground truth and estimated camera poses.
+Then, we calculate the $L_{2}$ distance of the ground truth and estimated reprojected 2D points.
+We use multiple thresholds, 1, 5, and 10 pixels, to evaluate the AUC scores.
 The camera translation and rotation error quantiles at 25%, 50%, and 75% are also reported.
-Besides, we also evaluate the matching quality by calculating the matching precision P, which is the ratio of inlier matches after PnP-RANSAC to the number of final matches $\mathcal{M}_{final}$.
+Moreover, we evaluate the matching quality by calculating the matching precision P, which is the ratio of inlier matches after PnP-RANSAC to the number of final matches $\mathcal{M}_{final}$.
 For visual localization tasks, we report the median translation (in meters) and rotation (in degrees) camera pose errors.
-Methods ScanNet dai2017scannet MegaDepth li2018megadepth
-Reproj. AUC (%) Rotation (${}^{\circ}$) Translation (m) P (%) ($\uparrow$) Reproj. AUC (%) Rotation (${}^{\circ}$) Translation P (%) ($\uparrow$)
-@1 / 5 / 10px ($\uparrow$) Quantile @25 / 50 / 75% ()($\downarrow$)( ↓ ) @1 / 5 / 10px ($\uparrow$) Quantile @25 / 50 / 75% ()($\downarrow$)( ↓ )
-k=1 Oracle 29.13 / 39.83 / 41.34 110.10 / 110.19 / 110.40 0.01 / 0.01 / 0.03 - 34.59 / 85.02 / 92.02 10.04 / 10.06 / 10.12 0.00 / 0.01 / 0.01 -
-BPnPNet campbell2020solving 10.00 / 10.00 / 10.02 199.17 / 128.90 / 154.68 4.35 / 6.82 / 9.86 13.60 10.22 / 10.63 / 10.89 16.13 / 32.01 / 61.58 1.67 / 3.17 / 5.44 12.95
-GoMatch zhou2022geometry 11.18 / 11.23 / 18.01 112.69 / 112.78 / 136.50 0.19 / 0.91 / 2.63 13.18 15.67 / 22.43 / 28.01 10.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73 14.94
-DGC-GNN 12.73 / 21.88 / 32.23 110.94 / 113.17 / 120.14 0.06 / 0.23 / 1.40 14.86 10.20 / 37.64 / 44.04 10.15 / 11.53 / 27.93 0.01 / 0.15 / 3.00 19.00
-k=10 BPnPNet campbell2020solving 10.00 / 10.00 / 10.03 104.68 / 135.94 / 160.54 4.67 / 7.30 / 10.92 10.84 10.36 / 10.72 / 10.97 16.63 / 34.69 / 67.77 1.64 / 3.30 / 5.97 10.74
-GoMatch zhou2022geometry 10.91 / 18.98 / 31.12 111.18 / 114.94 / 128.97 10.08 / 0.35 / 2.08 14.25 18.90 / 35.67 / 44.99 10.18 / 11.29 / 16.65 0.02 / 0.12 / 1.92 18.76
-DGC-GNN 11.76 / 31.74 / 48.11 110.67 / 111.49 / 117.62 10.04 / 0.11 / 0.53 16.42 15.30 / 51.70 / 60.01 10.07 / 10.26 / 15.41 0.01 / 0.02 / 0.57 13.36
-Table 1: 2D-3D matching.
-We report the reprojection error AUC scores thresholded at 1, 5, and 10 pixels; rotation (degrees) and translation error (meters) quantiles at 25, 50, and 75%; and matching precision (percentages).
-On MegaDepth li2018megadepth , the translation scale is unknown, so the unit of the errors comes from the reference reconstruction and is not in meters.
-Parameter $k$ is the number of images retrieved from the database to narrow down the search space.
-The best results are bold.
-DGC-GNN nearly doubles the AUC scores of GoMatch and reduces the pose errors to their ${\approx}33\%$.
 3.1 2D-3D Matching
-We compare with the two descriptor-free matchers GoMatch zhou2022geometry and BPnPnet campbell2020solving .
-During inference, we use the 3D points from the top-$k$ retrieved database images to match against the keypoints from query images.
-Following zhou2022geometry , we report the upper bound of the AUC score at 1, 5, and 10 pixels by the ground truth matches.
+We compare with the two descriptor-free matchers GoMatch [56] and BPnPnet [5].
+At inference, we use the 3D points from the top-$k$ retrieved database images to match with the keypoints from query images.
+Following [56], we report the upper bound of the AUC score using the ground truth matches.
 We refer to these values as Oracle.
-We use the official code with the default setting to generate the evaluation dataset on MegaDepth li2018megadepth and rerun the GoMatch and BPnPNet with the released models on the dataset.
+We select the GT matches by thresholding the reprojection error based on normalized image coordinates, using a threshold of $0.001$, to bypass the influence of camera intrinsics during GT selection.
+This is in contrast to what is done in [56].
+Results on GT selected by a pixel threshold are in the supp.
+material.
+We use the official code with the default setting to generate the evaluation dataset on MegaDepth [24] and rerun GoMatch and BPnPNet with the released models.
 Note that we also tested GoMatch after retraining it on MegaDepth and achieved similar results as with the released model.
+Methods G. Emb. C. Att. Color Ang. Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
+Sec. 2.2.2 Sec. 2.2.3 Sec. 2.2.1 Sec. 2.2.2 @1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
+GoMatch [56] 18.90 / 35.67 / 44.99 0.18 / 1.29 / 16.65 0.02 / 0.12 / 1.92
+Variants 10.86 / 41.18 / 50.51 0.13 / 0.76 / 13.47 0.01 / 0.07 / 1.62
+11.64 / 44.46 / 53.99 0.11 / 0.55 / 19.49 0.01 / 0.05 / 1.05
+13.20 / 46.33 / 54.34 0.09 / 0.41 / 19.98 0.01 / 0.03 / 1.19
+14.19 / 48.34 / 56.54 0.08 / 0.34 / 19.23 0.01 / 0.03 / 1.03
+DGC-GNN 15.30 / 51.70 / 60.01 0.07 / 0.26 / 15.41 0.01 / 0.02 / 0.57
+Table 2: Ablation Study.
+AUC scores thresholded at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, 75% with the proposed components added one by one to the GoMatch pipeline on the MegaDepth dataset.
 Matching Results.
 The results with $k=1$ and $k=10$ are presented in Table 1.
-$k$ is the number of retrieved image pairs that are used for evaluation.
+Parameter $k$ is the number of retrieved image pairs that are used for evaluation.
 The proposed method outperforms GoMatch and BPnPNet by a significant margin on both scenes.
@@
 Figure 4: Outlier Sensitivity.
-The AUC scores of BPnPNet campbell2020solving , GoMatch zhou2022geometry and the proposed DGC-GNN thresholded at 1, 5, and 10 pixels are plotted as a function of the outlier ratio.
-Oracle represents the AUC score upper bound using ground truth 2D-3D matches.
+The AUC scores of BPnPNet [5], GoMatch [56], and the proposed DGC-GNN thresholded at 1, 5, and 10 pixels are plotted as a function of the outlier ratio.
+Oracle represents the AUC upper bound using ground truth matches.
 Sensitivity to Outliers.
-To evaluate the sensitivity of our method to keypoint outliers, we follow the procedure in GoMatch zhou2022geometry .
+To evaluate the sensitivity to keypoint outliers, we follow the procedure in GoMatch [56].
 The outliers are controlled by the outlier ratio, ranging from 0 to 1, calculated as the number of unmatched keypoints divided by the maximum of the numbers of 2D and 3D points.
-In case the outlier ratio is 00, all of the input 2D and 3D points are selected from the ground truth matches, and no outliers are included in the matching process.
-When it is $1$, we directly use the keypoints from the 2D query and 3D points from top-$k$ retrieved images as input to find the matches without any filtering or outlier removal.
+If the outlier ratio is 00, all the input 2D and 3D points are selected from the ground truth matches, and no outliers are included in the matching process.
+When it is $1$, we directly use the keypoints from the query image and 3D points from the top-$k$ retrieved images without any filtering or outlier removal.
 The results are shown in Fig.
@@
 Even in the presence of outliers, DGC-GNN outperforms other methods by a large margin.
-This indicates that our method is more robust to the presence of outliers and can handle challenging matching scenarios more effectively than the state-of-the-art.
-Methods G. Emb. C. Att. Color Ang. Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
-Sec. 2.2.2 Sec. 2.2.3 Sec. 2.2.1 Sec. 2.2.2 @1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
-GoMatch zhou2022geometry 18.90 / 35.67 / 44.99 0.18 / 1.29 / 16.65 0.02 / 0.12 / 1.92
-Variants 10.86 / 41.18 / 50.51 0.13 / 0.76 / 13.47 0.01 / 0.07 / 1.62
-11.64 / 44.46 / 53.99 0.11 / 0.55 / 19.49 0.01 / 0.05 / 1.05
-14.19 / 48.34 / 56.54 0.08 / 0.34 / 19.23 0.01 / 0.03 / 1.03
-DGC-GNN 15.30 / 51.70 / 60.01 0.07 / 0.26 / 15.41 0.01 / 0.02 / 0.57
-Table 2: Ablation Study. AUC scores thresholded at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, 75% with the proposed components added one by one to the GoMatch pipeline.
+This indicates that our method is more robust to outliers and can handle challenging matching scenarios more effectively than the state-of-the-art.
 Ablation Study.
-We investigate the effectiveness of different components of DGC-GNN on the 2D-3D matching quality.
-The impact of different components of DGC-GNN is studied on MegaDepth dataset li2018megadepth with $k=10$.
+We investigate the effectiveness of different components of DGC-GNN on the 2D-3D matching quality on the MegaDepth dataset [24] with $k=10$.
 The results are reported in Table 2.
-We provide other results in the supplementary material.
+We provide results with $k=1$ in the supp.
+material.
 We conduct the ablations by gradually adding the components: global geometric embedding (G.
@@
 Incorporating color information into the matching process significantly impacts the performance, resulting in improvements of 2.55 / 3.88 / 2.55% (AUC@1 / 5 / 10px).
-This demonstrates the importance of considering color cues for accurate and robust matching.
-The global-to-local geometric embedding (G.
-Emb.) and Angular relation embedding (Ang.) substantially improve performance by 1.90 / 5.51 / 5.52% and 1.11 / 3.36 / 3.47%, respectively.
-It highlights the effectiveness of our idea incorporating global geometric context and local geometric details.
+This demonstrates the importance of color cues for accurate and robust matching.
+The global-to-local geometric (G.
+Emb.) and Angular relation embedding (Ang.) substantially improve the matching performance by 1.90 / 5.51 / 5.52% and 1.11 / 3.36 / 3.47%, respectively.
+It highlights the effectiveness of incorporating global geometric context and local geometric details.
 The cluster attention mechanism also plays a vital role, improving performance by 0.78 / 3.28 / 3.48%.
 The best results are obtained when all components are added to the pipeline.
-Methods No Desc. Maint. Privacy Cambridge-Landmarks kendall2015posenet (cm, ${}^{\circ}$) MB used 7Scenes shotton2013scene (cm, ${}^{\circ}$) MB used
+Methods No Desc. Maint. Privacy Cambridge-Landmarks [22] (cm, ${}^{\circ}$) MB used 7Scenes [46] (cm, ${}^{\circ}$) MB used
 King’s Hospital Shop St. Mary’s Chess Fire Heads Office Pumpkin Kitchen Stairs
-E2E MS-Trans. shavit2021learning 83 / 1.47 181 / 2.39 86 / 3.07 162 / 3.99 1171 11 / 4.66 24 / 9.60 14 / 12.19 17 / 5.66 18 / 4.44 17 / 5.94 26 / 8.45 11171
-DSAC* brachmann2021dsacstar 15 / 0.30 121 / 0.40 15 / 0.30 113 / 0.40 1112 12 / 1.10 12 / 1.24 11 / 1.82 13 / 1.15 14 / 1.34 14 / 1.68 13 / 1.16 11196
-HSCNet li2020hierarchical 18 / 0.30 119 / 0.30 16 / 0.30 119 / 0.30 1592 12 / 0.70 12 / 0.90 11 / 0.90 13 / 0.80 14 / 1.00 14 / 1.20 13 / 0.80 11036
-DB HybridSC camposeco2019hybrid 81 / 0.59 175 / 1.01 19 / 0.54 150 / 0.49 1113 - - - - - - - -
-AS sattler2016efficient 13 / 0.22 120 / 0.36 14 / 0.21 118 / 0.25 1813 13 / 0.87 12 / 1.01 11 / 0.82 14 / 1.15 17 / 1.69 15 / 1.72 4 / 1.01 -
-SP detone2018superpoint +SG sarlin2020superglue 12 / 0.20 115 / 0.30 14 / 0.20 117 / 0.21 3215 12 / 0.85 12 / 0.94 11 / 0.75 13 / 0.92 15 / 1.30 14 / 1.40 15 / 1.47 22977
-DF GoMatch zhou2022geometry 25 / 0.64 283 / 8.14 48 / 4.77 335 / 9.94 1148 14 / 1.65 13 / 3.86 19 / 5.17 11 / 2.48 16 / 3.32 13 / 2.84 89 / 21.12 11302
+E2E MS-Trans. [44] 83 / 1.47 181 / 2.39 86 / 3.07 162 / 3.99 1171 11 / 4.66 24 / 9.60 14 / 12.19 17 / 5.66 18 / 4.44 17 / 5.94 26 / 8.45 11171
+DSAC* [3] 15 / 0.30 121 / 0.40 15 / 0.30 113 / 0.40 1112 12 / 1.10 12 / 1.24 11 / 1.82 13 / 1.15 14 / 1.34 14 / 1.68 13 / 1.16 11196
+HSCNet [23] 18 / 0.30 119 / 0.30 16 / 0.30 119 / 0.30 1592 12 / 0.70 12 / 0.90 11 / 0.90 13 / 0.80 14 / 1.00 14 / 1.20 13 / 0.80 11036
+DB HybridSC [6] 81 / 0.59 175 / 1.01 19 / 0.54 150 / 0.49 1113 - - - - - - - -
+AS [41] 13 / 0.22 120 / 0.36 14 / 0.21 118 / 0.25 1813 13 / 0.87 12 / 1.01 11 / 0.82 14 / 1.15 17 / 1.69 15 / 1.72 4 / 1.01 -
+SP [11]+SG [38] 12 / 0.20 115 / 0.30 14 / 0.20 117 / 0.21 3215 12 / 0.85 12 / 0.94 11 / 0.75 13 / 0.92 15 / 1.30 14 / 1.40 15 / 1.47 22977
+DF GoMatch [56] 25 / 0.64 283 / 8.14 48 / 4.77 335 / 9.94 1148 14 / 1.65 13 / 3.86 19 / 5.17 11 / 2.48 16 / 3.32 13 / 2.84 89 / 21.12 11302
 DGC-GNN 18 / 0.47 75 / 2.83 15 / 1.57 106 / 4.03 1169 13 / 1.43 15 / 1.77 14 / 2.95 16 / 1.61 18 / 1.93 18 / 2.09 71 / 19.5 11355
 Table 3: Visual Localization.
-We report the median pose errors (cm, ${}^{\circ}$) and storage requirements (MB) on the scenes of the 7Scenes shotton2013scene and Cambridge-Landmarks kendall2015posenet datasets.
+We report the median pose errors (cm, ${}^{\circ}$) and storage requirements (MB) on the scenes of the 7Scenes [46] and Cambridge-Landmarks [22] datasets.
 Three groups of methods are shown: end-to-end (E2E), descriptor-based (DB), and descriptor-free (DF).
@@
 3.2 Visual Localization
-The visual localization task is the estimation of the 6 degrees-of-freedom camera pose of an input query image w.r.t a known map of the scene.
+Visual localization estimates the 6 degrees-of-freedom camera pose of an input query image w.r.t a known map of the scene.
 One of the most prominent ways of approaching this problem is via establishing 2D-3D correspondences and running robust pose estimation.
-Following zhou2022geometry , we ran the proposed DGC-GNN to obtain matches.
-For each query image, we match its keypoints against the 3D points from the top-10 retrieved views to build the 2D-3D correspondences.
-The camera pose is then estimated by PnP-RANSAC fischler1981random ; 1217599 .
-We consider two standard datasets 7scenes shotton2013scene and Cambridge landmarks kendall2015posenet to evaluate the proposed method.
-For 7scenes, we extract the keypoints with the SIFT detectors, and the top 10 pairs are retrieved using DenseVLAD torii201524 .
-For Cambridge landmarks, the keypoints are extracted by SuperPoint detone2018superpoint to ensure consistency with the SuperPoint-based structure-from-motion model.
-The top 10 pairs are provided by NetVLAD Arandjelovic16 .
-Results.
-In Table 3, we present the 3D model maintenance costs, privacy, storage requirements, and camera pose median errors (cm, ${}^{\circ}$) of standard descriptor-based localization techniques and descriptor-free methods, including the proposed DGC-GNN.
-The proposed method consistently outperforms GoMatch on all scenes by a significant margin.
-On the Cambridge-Landmarks dataset, the average median pose error of DGC-GNN is 54 cm / ${}^{\circ}$, while GoMatch leads to 173 cm / ${}^{\circ}$ average error.
-On 7Scenes, the average error of DGC-GNN is 15 cm / ${}^{\circ}$ and that of GoMatch is 22 cm / ${}^{\circ}$.
+Following [56], we ran the proposed DGC-GNN to obtain matches.
+For each query image, we match its keypoints with the 3D points from the top-10 retrieved views to build the 2D-3D correspondences.
+The camera pose is then estimated by PnP-RANSAC [16, 17].
+We use two standard datasets, 7Scenes [46] and Cambridge Landmarks [22].
+For 7Scenes, we extract the keypoints with the SIFT detectors, and the top 10 pairs are retrieved using DenseVLAD [50].
+For Cambridge Landmarks, the keypoints are extracted by SuperPoint [11] to ensure consistency with the SuperPoint-based structure-from-motion model.
+The top 10 pairs are provided by NetVLAD [1].
+3.2.1 Results
+In Table 3, we present the 3D model maintenance costs, privacy, storage requirements, and camera pose median errors (cm, ${}^{\circ}$) of standard descriptor-based localization techniques and descriptor-free methods.
+DGC-GNN consistently outperforms GoMatch on all scenes by a significant margin.
+On Cambridge Landmarks, the average median error of DGC-GNN is 54 cm / ${}^{\circ}$, while GoMatch leads to 173 cm / ${}^{\circ}$ error.
+On 7Scenes, the average error of DGC-GNN is 15 cm / ${}^{\circ}$, and that of GoMatch is 22 cm / ${}^{\circ}$.
 DGC-GNN requires a similar amount of memory to other descriptor-free methods.
-Also, it inherits their privacy-preserving properties due to not requiring and storing visual descriptors.
-The difference between descriptor-based (DB) and descriptor-free (DF) algorithms is clear.
+Also, it inherits their privacy-preserving properties due to not requiring visual descriptors.
+The trade-off between descriptor-based (DB) and descriptor-free (DF) algorithms is visible from the table.
 While descriptor-based ones lead to the best accuracy overall, they require excessive memory and descriptor maintenance and are susceptible to privacy attacks.
-Although the model compression method, HybridSC camposeco2019hybrid , shows effectiveness in storage saving, it achieves similar performance compared to DGC-GNN on the Cambridge dataset while still requiring descriptor maintenance.
+Although the model compression method, HybridSC [6], shows effectiveness in storage saving, it achieves similar performance compared to DGC-GNN on the Cambridge Landmarks dataset while still requiring descriptor maintenance.
 End-to-end methods (E2E) overcome these problems and achieve accurate results.
@@
 The proposed DGC-GNN only needs to be trained once, making it more efficient and convenient to use as an off-the-shelf tool.
+3.2.2 Generalizability
+Scenes Trained on MegaDepth [24] (SIFT) Trained on ScanNet [10] (SIFT)
+GoMatch (SIFT) GoMatch (SP) DGC-GNN (SIFT) DGC-GNN (SP) DGC-GNN (SIFT) DGC-GNN (SP)
+Chess 14 / 11.65 14 / 11.56 13 / 11.41 13 / 11.46 13 / 11.43 14 / 11.51
+Fire 13 / 13.86 12 / 13.71 15 / 11.81 17 / 12.30 15 / 11.77 16 / 12.03
+Heads 19 / 15.17 15 / 13.43 15 / 13.13 14 / 12.78 14 / 12.95 14 / 13.02
+Office 11 / 12.48 17 / 11.76 17 / 11.66 17 / 11.66 16 / 11.61 17 / 11.66
+Pumpkin 16 / 13.32 28 / 15.65 18 / 12.03 12 / 12.75 18 / 11.93 10 / 12.38
+Redkitchen 13 / 12.84 14 / 13.03 18 / 12.14 10 / 12.36 18 / 12.09 19 / 12.28
+Stairs 89 / 21.12 58 / 13.12 83 / 21.53 55 / 13.05 71 / 19.50 58 / 14.32
+All 22 / 15.78 18 / 14.61 17 / 14.82 14 / 13.77 15 / 14.47 14 / 13.89
+Table 4: Model Generalizability on Visual Localization Task.
+We report the translation and rotation median error (cm / ${}^{\circ}$) on 7Scenes dataset [46].
+We evaluate the models trained on MegaDepth [24] and ScanNet [10] datasets.
+The best performance is bold.
+Similar to [56], we discuss the generalizability of our DGC-GNN model on the visual localization task across different training and evaluation scenes.
+Specifically, we investigate the performance of our model when trained on MegaDepth [24] and ScanNet [10] and evaluated on the 7Scenes dataset [46].
+We also explore the impact of using different keypoint detectors, namely SIFT [27] and SuperPoint [11], during the evaluation.
+These experiments are summarized in Table 4, providing an overview of the performance of DGC-GNN under different training and evaluation conditions.
+While the best overall performance is achieved by training on the MegaDepth dataset with SIFT features, the results are similar in both training scenarios, showcasing that the proposed method generalizes well to unseen data.
+While we train on SIFT features, the best results are achieved by using SuperPoint features at inference time.
+This demonstrates that DGC-GNN is insensitive to the features used and can be utilized off-the-shelf even without retraining to our specific scenario.
 4 Conclusion
-In conclusion, this paper introduces DGC-GNN, a novel graph-based pipeline for descriptor-free 2D-3D matching that effectively leverages geometric and color cues in a global-to-local manner.
+In conclusion, this paper introduces DGC-GNN, a novel graph-based pipeline for visual descriptor-free 2D-3D matching that effectively leverages geometric and color cues in a global-to-local manner.
 Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming a geometric embedding to guide the local point matching.
 By employing a cluster-based transformer, we enable efficient information passing within local clusters, ultimately leading to significant improvements in the number of correct matches and the accuracy of pose estimation.
-Compared to the state-of-the-art descriptor-free matcher GoMatch zhou2022geometry , the proposed DGC-GNN demonstrates a substantial improvement, doubling the accuracy on real-world and large-scale datasets.
+Compared to the state-of-the-art descriptor-free matcher GoMatch [56], the proposed DGC-GNN demonstrates a substantial improvement, doubling the accuracy on real-world and large-scale datasets.
 Furthermore, it results in significantly increased localization accuracy.
 These advancements contribute to reducing the gap between descriptor-based and descriptor-free methods while addressing the limitations of descriptor-based ones, such as memory footprint, maintenance costs, and susceptibility to privacy attacks.
-The source code and models will be made publicly available.
 Limitations.
@@
 Acknowledgements.
-This work was supported by the Academy of Finland (grant No.
+This work was supported by the Academy of Finland (grants No.
 327911, No.
-353138), and by ETH Postdoc fellowship .
-We acknowledge the computational resources provided by the Aalto Science-IT project and CSC-IT Center for Science, Finland.
-5 Supplementary Material
+353138) and the Hasler Stiftung Research Grant via the ETH Zurich Foundation.
+We acknowledge the computational resources provided by the CSC-IT Center for Science, Finland.
+Supplementary Material
 Appendix A Training and Evaluation Details
 Dataset Generation.
-The training data generation process for MegaDepth li2018megadepth follows the methodology outlined in GoMatch zhou2022geometry .
-The undistorted SfM model reconstructions used in MegaDepth are provided by D2Net dusmanu2019d2 .
+The training data generation process for MegaDepth [24] follows the methodology outlined in GoMatch [56].
+The undistorted SfM model reconstructions used in MegaDepth are provided by D2Net [14].
 For training, we sample up to 500 images from each scene.
 For each sampled image, we select the top-$k$ co-visible views that have at least 35% image overlap.
-This ensures that there are enough matches available for training.
+This ensures that there are enough matches for training.
 The overlapping score is computed by dividing the number of co-visible 3D points by the total number of points in the training image.
-In the case of ScanNet dai2017scannet , a similar procedure is conducted.
+In the case of ScanNet [10], a similar procedure is conducted.
 We also sample up to 500 images from each scene for the training set generation.
-The co-visible images are obtained using the co-visible scores provided by LOFTR sun2021loftr .
-We extract all the co-visible views of the training image with co-visible scores larger than 0.65, and then randomly sample the top-$k$ views for training.
-Since ScanNet is an RGB-D dataset without SfM models, we obtain the 3D points for each image by projecting the detected 2D keypoints with valid depth to the 3D coordinate.
+The co-visible images are obtained using the co-visible scores provided by LoFTR [49].
+We extract all the co-visible views of a training image with co-visible scores larger than 0.65.
+Then, we randomly sample the top-$k$ views for training.
+Since ScanNet is an RGB-D dataset without an SfM reconstruction, we obtain the 3D points for each image by projecting the detected 2D keypoints with valid depth to 3D.
 By doing this for each image, we reconstruct a sparse 3D point cloud based on the detected 2D keypoints.
@@
 The test set for ScanNet consists of 14,892 query images from 30 scenes.
-The data generation of 7Scenes shotton2013scene and Cambridge Landmarks kendall2015posenet follows the same procedure in  zhou2022geometry .
+The data generation of 7Scenes [46] and Cambridge dataset [22] follows the same procedure in  [56].
+Figure 5: Points Reprojection and Image Recovery Example.
+Method Reprojection InvSFM [35]
+Points Points+RGB Points Points+RGB Points+SIFT
+SSIM ($\downarrow$) 0.240 0.258 0.352 0.375 0.476
+Table 5: SSIM Results. We evaluate the SSIM from Point Reprojection and Image Recovering, adding RGB to points leads only to a slight SSIM increase on both reprojection and image recovery.
 Inference.
@@
 The 3D points from the top-$k$ retrieved database images are then applied to match against the queries with our proposed pipeline.
-We use the Sinkhorn algorithm sinkhorn1967concerning ; cuturi2013sinkhorn to optimize the extended cost matrix ¯$\bar{\mathcal{M}}\in\mathbb{R}^{N+1,M+1}$ in an iterative manner with up to 20 iterations to obtain the initial matches.
+We use the Sinkhorn algorithm [47, 9] to optimize the extended cost matrix ¯$\bar{\mathcal{M}}\in\mathbb{R}^{N+1,M+1}$ in an iterative manner with up to 20 iterations to obtain the initial matches.
 The final matches are obtained by filtering the matches with matching confidence $\theta<0.5$ in the outlier rejection module.
-For the visual localization task, the camera poses are estimated by the P3P solver with RANSAC fischler1981random implemented in OpenCV bradski2000opencv and then refined by Levenberg-Marquardt more2006levenberg algorithm on the inliers matches, minimising the reprojection error.
-Figure 5: 2D-3D Matching (shown by green lines) with the proposed DGC-GNN and GoMatch zhou2022geometry .
-Appendix B Additional Results
+For the visual localization task, the camera poses are estimated by the P3P solver with RANSAC [16] implemented in OpenCV [4] and then refined by Levenberg-Marquardt [29] algorithm on the inliers matches, minimizing the reprojection error.
+Figure 6: 2D-3D Matching (shown by green lines) with the proposed DGC-GNN and GoMatch [56].
+Appendix B Privacy Issue of RGB Points
+We investigate the impact on privacy resulting from incorporating RGB information into pixels and points.
+To assess this, we compute the Structural Similarity Index Measure (SSIM) for 3D points reprojected onto the image plane against the ground truth (GT) images on MegaDepth over 500 images from multiple scenes.
+Additionally, we recover the images from points + RGB and points + descriptors with InvSFM [35] to calculate the SSIM against the GTs.
+The findings are detailed in Table 5 and Fig 5.
+The addition of RGB data to points results in only a marginal increase in SSIM for both direct reprojection and image reconstruction via InvSFM, significantly less than what is achieved by incorporating SIFT descriptors.
+It is worth noting that denser point clouds might provide sufficient context, potentially leading to privacy concerns.
+However, in our setting, we mitigate this risk by limiting the number of keypoints from each database image to a maximum of 1024.
+Appendix C Additional Results
 Qualitative Results.
 More visualizations of inlier matches provided by DGC-GNN and GoMatch on MegaDepth are shown in Fig.
-5.
-DGC-GNN consistently achieves more matches on multiple scenes, highlighting the effectiveness of the proposed method.
+6.
+DGC-GNN consistently finds more correct matches on multiple scenes, highlighting the effectiveness of the proposed method.
 Methods Global C. Att. Color Ang. Cluster Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
 @1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
-GoMatch zhou2022geometry 15.67 / 22.43 / 28.01 0.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73
+GoMatch [56] (w/o OR) 14.47 / 17.95 / 23.42 1.29 / 11.85 / 33.60 0.11 / 1.18 / 3.58
+GoMatch [56] 15.67 / 22.43 / 28.01 0.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73
 Variants G.Emb K-means 17.68 / 28.41 / 34.36 0.28 / 16.78 / 34.52 0.03 / 0.73 / 3.77
@@
 G.Emb Mean-shift 10.07 / 36.01 / 43.03 0.16 / 12.15 / 28.99 0.01 / 0.20 / 3.26
+DGC-GNN (w/o OR) G.Emb K-means 18.56 / 30.79 / 37.03 0.22 / 14.85 / 30.07 0.02 / 0.47 / 3.10
 DGC-GNN G.Emb K-means 10.20 / 37.64 / 44.04 0.15 / 11.53 / 27.93 0.01 / 0.15 / 3.00
@@
 Figure 7: Qualitative Matching Results of Different Architectures. We visualize the number of inlier matches after the PnP-RANSAC with different architectures (shown by green lines ).
-Figure 7: Architectural Ablation Study.
-We plot the training loss and the number of matches estimated during the training for different architectures.
-Notably, we observe that the inclusion of color encoding in the training procedure leads to a significant improvement.
 Additional Ablation Results.
-In addition to the ablation results presented in the main paper, we also provide ablation results for single-view matching with $k=1$ on MegaDepth li2018megadepth .
+In addition to the ablation results presented in the main paper, we also provide ablation results for single-view matching with $k=1$ on MegaDepth [24].
 Furthermore, we conduct two additional ablations to investigate the impact of different component selections.
@@
 We compare the performance of K-Means and Mean-Shift clustering algorithms in our pipeline.
+Last, we study the effectiveness of the outlier rejection (OR) network.
 The results are presented in Table 6.
@@
 The results obtained using the global label embedding (G.
-Label) with cluster attention (C.Att).
-show even worse performance compare to geometric embedding (G.
+Label) with cluster attention (C.Att) show even worse performance compared to geometric embedding (G.
 Emb.) only, indicating the superiority of our clustering-based geometric embedding over the label embedding and highlighting the importance of incorporating geometric cues in the embedding process for effective point matching.
 Regarding the impact of different clustering algorithms, we only observe a minor difference in K-Means and Mean-Shift results, suggesting that our approach is robust to the choice of the clustering algorithm.
-In addition to the numerical results, we also visualize the inlier matches (see Fig.
-6), the training loss and the number of predicted matches during training (see Fig.
-7) with different architectures to provide deep insights into the behaviour and performance of different architectures.
-Appendix C Generalizability
-Similar to zhou2022geometry , we discuss the generalizability of our DGC-GNN model on the visual localization task across different types of training and evaluation scenes.
-Specifically, we investigate the performance of our model when trained on MegaDepth li2018megadepth and ScanNet dai2017scannet , and evaluated on the 7Scenes dataset shotton2013scene .
-We also explore the impact of using different keypoint detectors, namely SIFT lowe2004distinctive and SuperPoint detone2018superpoint , during the evaluation.
-The results of these experiments are summarized in Table 5, which provides an overview of the performance of our DGC-GNN model under different training and evaluation conditions.
-The results show that the DGC-GNN model trained on the ScanNet dataset, using the SIFT keypoint detector, and evaluated on the indoor 7Scenes dataset with the same detector, achieves the best overall performance.
-This demonstrates the effectiveness and robustness of our model in this specific training and evaluation scenario.
-However, we also note minor differences in performance when the model is trained on the MegaDepth dataset or evaluated using SuperPoint keypoints.
-Despite these variations, our findings suggest that the DGC-GNN model is scene-agnostic and can generalize well to different types of keypoint detectors.
-This is an encouraging result as showcases the versatility and generalization capability of the DGC-GNN approach, allowing it to be applied in various settings without significant degradation in performance.
-Scenes Trained on MegaDepth li2018megadepth (SIFT) Trained on ScanNet dai2017scannet (SIFT)
-GoMatch zhou2022geometry (SIFT) GoMatch zhou2022geometry (SP) DGC-GNN (SIFT) DGC-GNN (SP) DGC-GNN (SIFT) DGC-GNN (SP)
-Chess 14 / 11.65 14 / 11.56 13 / 11.41 13 / 11.46 13 / 11.43 14 / 11.51
-Fire 13 / 13.86 12 / 13.71 15 / 11.81 17 / 12.30 15 / 11.77 16 / 12.03
-Heads 19 / 15.17 15 / 13.43 15 / 13.13 14 / 12.78 14 / 12.95 14 / 13.02
-Office 11 / 12.48 17 / 11.76 17 / 11.66 17 / 11.66 16 / 11.61 17 / 11.66
-Pumpkin 16 / 13.32 28 / 15.65 18 / 12.03 12 / 12.75 18 / 11.93 10 / 12.38
-Redkitchen 13 / 12.84 14 / 13.03 18 / 12.14 10 / 12.36 18 / 12.09 19 / 12.28
-Stairs 89 / 21.12 58 / 13.12 83 / 21.53 55 / 13.05 71 / 19.50 58 / 14.32
-Table 5: Model Generalizability on Visual Localization Task.
-We report the translation and rotation median error (cm / ${}^{\circ}$) on 7Scenes dataset shotton2013scene .
-We evaluate the model trained on both MegaDepth li2018megadepth and ScanNet dai2017scannet datasets.
-The best performance is bold.
+The results also demonstrate that outlier rejection is an essential post-processing module to achieve good performance.
+In addition to the numerical results, we visualize the inlier matches (see Fig.
+7) to provide deep insights into the behavior and performance of different architectures.
+Hyperparameters analysis.
+Besides the component ablations, we also give an in-depth analysis of the hyperparameters used in our main pipeline.
+Here, we add additional ablations on the number of input keypoints, the number of cluster groups at the coarse level, the number of nearest neighbors in the local graph build, and the outlier rejection threshold by retraining our DGC-GNN.
+The results are presented in Table 7.
+We observe that DGC-GNN with G.
+Clusters = 10 and Local NN = 10 achieves overall the best performance.
+Setting the outlier rejection threshold to 0.7 leads to the best performance.
+However, the results are stable across different configurations, indicating robustness to the parameter setting.
+Methods G. Cluters Local NN OR Threshold Reproj. AUC (%) Rotation (${}^{\circ}$) Translation
+@1 / 5 / 10px ($\uparrow$) Quantile@25 / 50 / 75% ($\downarrow$)
+DGC-GNN 10 10 0.5 15.30 / 51.70 / 60.01 0.07 / 0.26 / 15.41 0.01 / 0.02 / 0.57
+HyperParam. 5 10 0.5 14.73 / 50.12 / 58.26 0.08 / 0.28 / 18.76 0.01 / 0.03 / 0.99
+15 10 0.5 15.14 / 50.56 / 58.62 0.07 / 0.28 / 17.66 0.01 / 0.03 / 0.89
+10 20 0.5 14.77 / 49.84 / 57.97 0.07 / 0.29 / 18.26 0.01 / 0.03 / 0.90
+10 30 0.5 14.75 / 50.95 / 59.45 0.08 / 0.28 / 15.48 0.01 / 0.03 / 0.58
+10 10 0.3 13.28 / 46.44 / 55.05 0.08 / 0.43 / 18.63 0.01 / 0.04 / 0.98
+10 10 0.7 16.63 / 56.26 / 64.46 0.07 / 0.19 / 12.58 0.01 / 0.02 / 0.27
+Table 7: Ablation Study on Hyperparameters.
+We report the results of ablations with retrieved image $k=10$ on the number of global clusters, the number of nearest neighbour points for local graph build, and different thresholds for outlier rejection.
+The best results are bold.
+Matching Results in pixel threshold.
+As mentioned in the main paper, we selected the ground truth matches in normalized image coordinates.
+The described GT difference only affects the reprojection AUC scores.
+Here, we present the matching results in Table 5 by selecting the ground truth matches in pixel coordinate with $1$ pixel threshold as done in [56].
+Our conclusions still hold.
 Appendix D Model Parameters and Timing
 We discuss the model parameters and running time of DGC-GNN in this section.
-DGC-GNN incorporate global geometric embedding and local clustering attention, which has around 5.7 million trainable parameters and an estimated model size of 22.6 MB.
-The average inference time for each image pair over the Megadepth evaluation queries is 77.8ms, and the number is 70.9ms on ScanNet queries.
-The measurements are conducted on a 32GB NVIDIA Telsa V100 GPU.
+DGC-GNN incorporates global geometric embedding and local clustering attention, which has around 5.7 million trainable parameters and an estimated model size of 22.6 MB.
+The average inference time for each image pair over the Megadepth evaluation queries is 77.8ms.
+It roughly breaks down into point encoding (24 ms), global geometric embedding (14 ms), cluster-based attention (22 ms), optimal transport (7 ms), and outlier rejection (8 ms).
+The measurements are conducted on a 32GB NVIDIA Telsa V100 GPU with a maximum of 1024 keypoints.